### v1

# An El-Zahar Type Theorem in $3$-graphs under Codegree Condition

Abstract
A $3$-uniform loose cycle, denoted by $C_{t}$, is a $t$-vertex 3-uniform hypergraph whose vertices can be ordered cyclically in such a way that the hyperedges are sets of consecutive three vertices and every two consecutive hyperedges share exactly one vertex. The length of $C_{t}$ is the number of its edges. We prove that for any $\eta>0$, there exists an $n_{0}=n_{0}(\eta)$ such that for any $n\geq n_{0}$ the following holds. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$ such that $\sum_{i=1}^{r}n_{i}=n$ with each loose cycle having a length of at least $3$. Let $k$ be the number of loose cycles with odd lengths in $\mathcal{C}$. If $\mathcal{H}$ is a $3$-uniform hypergraph on $n$ vertices with minimum codegree at least $(n+2k)/4+\eta n$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph. This generalizes the results of Kühn and Osthus for loose Hamilton cycle and the result of Mycroft for loose cycle factors in $3$-graphs, where each cycle has length no less than $3$. The codegree bound is approximately tight. Our proof relies on a transversal blow-up lemma recently developed by the first author and Staden.
Keywords: loose cycles, codegree, regularity lemma, transversal blow-up lemma
1 Introduction
In 1952, Dirac [10] proved that every $n$-vertex graph with a minimum degree of at least $n/2$ contains a Hamilton cycle. This has motivated extensive research into determining the minimum degree required to ensure a spanning subgraph in a given graph $G$, i.e., a subgraph that has as many vertices as $G$. Determining the minimum degree for vertex-disjoint copies of a given graph has garnered significant attention. Given a graph $G$ of order $n$ and a graph $F$ of order $h$, an $F$-tiling of $G$ is a subgraph of $G$ that consists of vertex-disjoint copies of $F$. We say that $G$ has a perfect $F$-tiling (or an $F$-factor) if $h$ divides $n$ and $G$ contains $n/h$ vertex-disjoint copies of $F$. Hell and Kirkpatrick [17] showed that when the graph $F$ has any component with at least three vertices, the existing question of whether a graph $G$ has an $F$-factor is NP-complete.
This leads to a natural problem: finding sufficient conditions for $G$ to have an $F$-factor. The celebrated Hajnal-Szemerédi [14] theorem shows that every $n$-vertex graph $G$ with minimum degree at least $(k-1)n/k$ must contain a $K_{k}$-factor. The case $k=3$ was proved by Corrádi and Hajnal [6]. For any fixed graph $F$, Alon and Yuster [3] initially characterized the minimum degree of $G$ that ensures an $F$-factor by using the chromatic number of $F$. Kühn and Osthus [20] finally determined the exact minimum degree of a graph $G$ that ensures an $F$-factor.
For graphs, a classical conjecture of El-Zahar states the following [11].
Conjecture 1.1 ([11], El-Zahar’s conjecture).
Let $H$ be a graph consisting of $r$ vertex-disjoint cycles of order $n_{1},n_{2},\ldots,n_{r}$ satisfying $n_{1}+n_{2}+\cdots+n_{r}=n$, and $G$ be a graph on $n$ vertices with minimum degree at least $\sum_{i=1}^{r}\lceil n_{i}/2\rceil$, then $G$ contains $H$ as a subgraph.
Abbasi [1] confirmed this conjecture for sufficiently large $n$. Note that when $r=1$, it is Dirac’s theorem. The famous Corrádi-Hajnal theorem [6] follows from the special case $n_{1}=n_{2}=\ldots=n_{r}=3$, i.e., triangles. The case when $n_{1}=n_{2}=\ldots=n_{i}=4$ is an old conjecture of Erdős and Faudree [12], which was solved by Wang [30].
Over the past few decades, there has been substantial interest in extending the tiling problems into hypergraphs. However, the problems become much harder for hypergraphs. Given an integer $k\geq 2$, a $k$-uniform hypergraph, often referred to as a $k$-graph, consists of a vertex set $V$ and an edge set $E\subseteq\binom{V}{k}$.
For a $3$-graph $\mathcal{H}=(V,E)$ and two distinct vertices $u,v\in V$, let $N_{\mathcal{H}}(v):=\{\{x,y\}\in\binom{V}{2}:\{v,x,y\}\in E(\mathcal{H})\}$ and $N_{\mathcal{H}}(u,v):=\{w\in V:\{u,v,w\}\in E(\mathcal{H})\}$. The degree of a vertex $v\in V$, denoted by $d_{\mathcal{H}}(v)$, is defined to be $|N_{\mathcal{H}}(v)|$. The minimum vertex degree of $\mathcal{H}$ is defined as $\delta_{1}(\mathcal{H}):=\min_{v\in V}|N_{\mathcal{\mathcal{H}}}(v)|$, and the minimum codegree of $\mathcal{H}$ is defined as $\delta_{2}(\mathcal{H}):=\min_{u,v\in\binom{V}{2}}|N_{\mathcal{\mathcal{H}}}(u,v)|$.
A $t$-vertex $3$-graph is called a loose path, denoted by $P_{t}$, if its vertices can be ordered as $v_{1},v_{2},\ldots,v_{t}$ so that $E(P_{t})=\{\{v_{2i+1},v_{2i+2},v_{2i+3}\}:i=0,\ldots,(t-3)/2\}$. Note that $t$ must be odd. Once the order of $P_{t}$ is determined, we call a vertex that lies in its first hyperedge but not in the second one the endpoints of $P_{t}$. A $t$-vertex $3$-graph is called a loose cycle, denoted by $C_{t}$, if its vertices can be ordered as $v_{1},v_{2},\ldots,v_{t}$ so that
$E(C_{t})=\{\{v_{2i+1},v_{2i+2},v_{2i+3}\}:i=0,\ldots,(t-4)/2\}\cup\{v_{1},v_{t},v_{t-1}\}.$
Note that $t$ must be even. The length of $P_{t}$ (or $C_{t}$) is the number of its hyperedges, denoted by $\ell(P_{t})$ (or $\ell(C_{t})$) in the following. By the definitions, we know that $\ell(P_{t})=(t-1)/2$ and $\ell(C_{t})=t/2$.
For any two $3$-graphs $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, an embedding from $\mathcal{H}_{1}$ into $\mathcal{H}_{2}$ is an injective map from $V(\mathcal{H}_{1})$ to $V(\mathcal{H}_{2})$ that preserves adjacency. Moreover, if $|V(\mathcal{H}_{1})|=|V(\mathcal{H}_{2})|$, we say the embedding is a perfect embedding.
In hypergraphs, there are various notions of minimum degree and cycles for $k$-graphs. This leads to a few variants of the same meta-problem: which minimum degree conditions on a $k$-graph $\mathcal{H}$ are sufficient for the existence of an $F$-tiling for a $k$-graph $F$? In this paper, we focus on determining the minimum codegree in $3$-graphs that ensures spanning disjoint loose cycles.
As a natural generalization of Dirac’s theorem, Kühn and Osthus [21] gave an approximate bound that ensures a loose Hamilton cycle in a $3$-graph.
Theorem 1.2 ([21]).
For each $\sigma>0$, there is an integer $n_{0}=n_{0}(\sigma)$ such that every $3$-graph $\mathcal{H}$ with $n\geq n_{0}$ vertices and minimum codegree at least $(1/4+\sigma)n$ contains a loose Hamilton cycle.
Czygrinow and Molla [9] determined that the minimum codegree of $n/4$ suffices. In [21], Kühn and Osthus also showed that every $n$-vertex $3$-graph with minimum codegree $\delta_{2}(\mathcal{H})\geq(1/4+o(1))n$ contains a $C_{4}$-factor, and Czygrinow, DeBiasio and Nagle [7] determined the codegree exactly for large $n$. Gao and Han [13] determined that the tight minimum codegree that forces a $C_{6}$-factor in an $n$-vertex $3$-graph is $n/3$. To answer a question of Rödl and Ruciński ([25], Problem 3.15) and to extend the results to a more general case, Czygrinow [8] established the codegree condition for a $C_{s}$-factor in a $3$-graph for every even integer $s\geq 6$:
Theorem 1.3 ([8]).
For every even integer $s\geq 6$, there is $n_{0}$ such that every $3$-graph $\mathcal{H}$ on $n\in s\mathbb{Z},n\geq n_{0}$ vertices with $\delta_{2}(H)\geq\frac{\lceil s/4\rceil}{s}n$, has a $C_{s}$-factor.
This bound is tight. Mycroft [24] established an asymptotic minimum codegree condition for loose cycle factor in $k$-graphs. When $k=3$, up to an $o(n)$ error term, this bound meets Czygrinow’s bound for $C_{s}$-factor with $s$ even.
In this paper we consider an analogue of El-Zahar’s conjecture for $3$-graphs, and we have the following result:
Theorem 1.4 (Main result).
For every $\eta>0$, there exists an integer $n_{0}=n_{0}(\eta)$ such that the following holds for any $n\geq n_{0}$. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$, where $\sum_{i=1}^{r}n_{i}=n$, and each loose cycle has a length of at least 3. Let $k$ denote the number of loose cycles with odd lengths. If $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq(n+2k)/4+\eta n$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph.
Note that the case when $k=0$ means that there is no loose cycle with odd length in $\mathcal{C}$. We remark that the two results of Kühn and Osthus [21] follow from the cases when $r=1$ (Hamilton cycle) and $n_{1}=n_{2}=\cdots=n_{r}=4$. Theorem 1.3 follows from the case when $n_{1}=n_{2}=\cdots=n_{r}=n/s$. The bound on the minimum codegree in Theorem 1.4 is asymptotically tight. Indeed, if $n_{1}=n_{2}=\cdots=n_{r}=6$, that is, the $3$-graph $\mathcal{C}$ contains $n/6$ loose cycles $C_{6}$, then $\delta_{2}(\mathcal{H})\geq(n+2k)/4+\eta n=(1/3+\eta)n$. This implies that the bound is asymptotically tight by the result of Mycroft in [24].
2 Sketch of the Proof
For 2-graphs, the ‘regularity-blow-up method’ has been applied to solve many spanning subgraph embedding problems, see [18, 19, 23] and the references. The proof of Theorem 1.4 uses the weak regularity method and blow-up lemma for transversals. We first reduce the problem to the case where the length of each cycle is at least 6 and at most $\log n$, and we divide the proof into the following three steps:
Proof Steps:
Step 1: Apply the regularity method to obtain the reduced hypergraph and find an almost perfect tiling of some defined $3$-graph. Next, we recover each component of this tiling back to $\mathcal{H}$, and perform a random slicing to obtain some regular triples $T(X,Y,Z)$ that cover almost all of the vertex set, where the sizes of $X$, $Y$, and $Z$ satisfy a specific proportional relationship. Using the slicing lemma, we can ensure that these triples are half-superregular.
Step 2: Embed some cycles from the list to use up the vertices not contained in any of the triples. Embed more cycles in the list to consume some of the vertices in the triples such that each tripe has only been consumed a little fraction of vertices, and the remaining cycles to be embedded could be partitioned into pieces that allows us to build an injection from the piece set to the triple set.
Step 3: For each triple, apply the transversal blow-up lemma to embed the set of cycles from the corresponding piece.
3 Preliminaries
In this section, we present some useful lemmas that we will use in the next section, involving the weak hypergraph regularity lemma, the transversal blow-up lemma, and probabilistic tools.
3.1 Regularity method for 3-graphs
In this subsection, we review the well-known weak hypergraph regularity lemma, a straightforward extension of Szemerédi’s regularity lemma for graphs [28]. Before our paper, the strong regularity lemma was commonly used in the spanning subgraph embedding problem for hypergraphs, see [15, 21, 26, 27] and the references. In our paper, we only need to use the weak regularity lemma.
Let $\mathcal{H}$ be a $3$-partite $3$-graph with disjoint vertex sets $V_{1},V_{2}$, and $V_{3}$, which we also denote as $(V_{1},V_{2},V_{3})_{\mathcal{H}}$. Each hyperedge in $\mathcal{H}$ contains exactly one vertex from each set $V_{i}$ for $i\in[3]$. Let $e(\mathcal{H})$ denote the number of edges in $\mathcal{H}$. The density of $\mathcal{H}$ is defined as:
$d_{\mathcal{H}}(V_{1},V_{2},V_{3})=\frac{e(\mathcal{H})}{|V_{1}||V_{2}||V_{3}|}.$
Given $\varepsilon>0$, we say that $(V_{1},V_{2},V_{3})_{\mathcal{H}}$ is:
  • $\varepsilon$-regular if for every sub-hypergraph $(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})_{\mathcal{H}}$ with $V_{i}^{\prime}\subseteq V_{i}$ and $|V_{i}^{\prime}|\geq\varepsilon|V_{i}|$ for each $i\in[3]$, we have
    $|d_{\mathcal{H}}(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})-d_{\mathcal{H}}(V_{1},V_{2},V_{3})|<\varepsilon;$
  • $(\varepsilon,d)$-regular if additionally $d_{\mathcal{H}}(V_{1},V_{2},V_{3})\geq d$;
  • $(\varepsilon,d)$-half-regular if for every sub-hypergraph $(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})_{\mathcal{H}}$ with $V_{i}^{\prime}\subseteq V_{i}$ and $|V_{i}^{\prime}|\geq\varepsilon|V_{i}|$ for each $i\in[3],$ we have $d_{\mathcal{H}}(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})\geq d$;
  • $(\varepsilon,d)$-superregular if it is $(\varepsilon,d)$-regular and additionally,
    $d_{\mathcal{H}}(x)\geq\frac{d|V_{1}||V_{2}||V_{3}|}{|V_{i}|}$
    for each $i\in[3]$ and $x\in V_{i}$;
  • $(\varepsilon,d)$-half-superregular if for every sub-hypergraph $(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})_{\mathcal{H}}$ with $V_{i}^{\prime}\subseteq V_{i}$ and $|V_{i}^{\prime}|\geq\varepsilon|V_{i}|$ for each $i\in[3],$ we have $d_{\mathcal{H}}(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})\geq d$, and
    $d_{\mathcal{H}}(x)\geq\frac{d|V_{1}||V_{2}||V_{3}|}{|V_{i}|}$
    for each $i\in[3]$ and $x\in V_{i}$.
The following so-called slicing lemma, introduced in [4], where the case of “half-(super)regular” can be obtained based on the original proof with some additional trivial analysis, is a useful lemma in this paper. For any two constants $a,b\in(0,1)$, we write $a\ll b$ if there exists a function $a_{0}=a_{0}(b)$ such that the subsequent arguments hold for all $0<a\leq a_{0}$. When we write multiple constants in a hierarchy, we mean they are chosen from right to left.
Lemma 3.1 ([4], Slicing lemma).
Let $0<1/m\ll\varepsilon\ll\alpha\ll d\leq 1$. Let $\mathcal{H}$ be a $3$-partite $3$-graph with vertex sets $V_{1},V_{2},V_{3}$ and $m\leq\min\{|V_{1}|,|V_{2}|,|V_{3}|\}$. Let $V^{\prime}_{i}\subseteq V_{i}$ for each $i\in[3]$ and $\mathcal{H^{\prime}}$ be the sub-hypergraph of $\mathcal{H}$ induced on $V_{1}^{\prime}\cup V_{2}^{\prime}\cup V_{3}^{\prime}$. Then the following holds.
  1. $(i)$
    Suppose that $\mathcal{H}$ is $(\varepsilon,d)$-(half)-regular, and $|V^{\prime}_{i}|\geq\alpha|V_{i}|$ for each $i\in[3]$. Then $\mathcal{H^{\prime}}$ is $(\varepsilon/\alpha,d/2)$-(half)-regular.
  2. $(ii)$
    Suppose that $\mathcal{H}$ is $(\varepsilon,d)$-(half)-superregular, and $|V^{\prime}_{i}|\geq(1-\alpha)|V_{i}|$ for each $i\in[3]$. Then $\mathcal{H^{\prime}}$ is $(2\varepsilon,d/2)$-(half)-superregular.
  3. $(iii)$
    Suppose that $\mathcal{H}$ is $(\varepsilon,d)$-(half)-superregular, and $m_{i}\geq\alpha|V_{i}|$ for each $i\in[3]$. If $V^{\prime}_{i}\subseteq V_{i}$ is a uniform random subset of size $m_{i}$ for each $i\in[3]$, then with high probability, $\mathcal{H^{\prime}}$ is $(\varepsilon/\alpha,d^{2}/16)$-(half)-superregular.
Using the Slicing lemma, it is easy to obtain the following facts. It shows that we can extract a superregular $3$-partite sub-hypergraph from a given regular $3$-partite $3$-graph by removing a small number of vertices from each vertex set.
Lemma 3.2.
Let $0<\varepsilon\ll d\leq 1$, and let $\mathcal{H}$ be an $(\varepsilon,d)$-(half)-regular $3$-partite $3$-graph with vertex sets $V_{1},V_{2}$, and $V_{3}$. Then there exist subsets $V_{i}^{\prime}\subseteq V_{i}$ with $|V_{i}^{\prime}|=(1-\varepsilon)|V_{i}|$ for each $i\in[3]$ such that $(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})_{\mathcal{H}}$ is $(2\varepsilon,d/2)$-(half)-superregular.
Proof.
Let $V_{i}^{\prime\prime}:=\{v\in V_{i}:d_{\mathcal{H}}(v)<(d-3\varepsilon)|V_{j}||V_{h}|\}$ for $\{j,h\}=[3]\setminus\{i\}$ and each $i\in[3]$. We claim that $|V_{i}^{\prime\prime}|<\varepsilon|V_{i}|$ for each $i\in[3]$. Indeed, suppose for contradiction that $|V_{1}^{\prime\prime}|\geq\varepsilon|V_{1}|$. Then we have $d_{\mathcal{H}}(V_{1}^{\prime\prime},V_{2},V_{3})>d-\varepsilon$. However, by the definition of $V_{1}^{\prime\prime}$, we have $e_{\mathcal{H}}(V_{1}^{\prime\prime},V_{2},V_{3})<(d-3\varepsilon)|V_{1}^{\prime\prime}||V_{2}||V_{3}|$. This implies that $d_{\mathcal{H}}(V_{1}^{\prime\prime},V_{2},V_{3})<d-3\varepsilon$, a contradiction.
Let $V_{i}^{\prime}\subseteq V_{i}\setminus V_{i}^{\prime\prime}$ such that $|V_{i}^{\prime}|=(1-\varepsilon)|V_{i}|$ for each $i\in[3]$. We show that $\mathcal{H}^{\prime}=(V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime})_{\mathcal{H}}$ is $(2\varepsilon,d-3\varepsilon)$-(half)-superregular. By Lemma 3.1 (i), $\mathcal{H}^{\prime}$ is $(2\varepsilon,d/2)$-(half)-regular. Moreover, for each vertex $v\in V_{i}^{\prime}$ and $\{j,h\}=[3]\setminus\{i\}$, we have
$d_{\mathcal{H}^{\prime}}(v)\geq(d-3\varepsilon)|V_{j}||V_{h}|-2\varepsilon|V_{i}||V_{j}|\geq d|V^{\prime}_{j}||V^{\prime}_{h}|/2.$
Therefore, $\mathcal{H^{\prime}}$ is $(2\varepsilon,d/2)$-(half)-superregular. ∎
Now, we present the weak hypergraph regularity lemma, which follows from Chung [5].
Lemma 3.3 (Weak hypergraph regularity lemma).
For all integers $L_{0}\geq 1$ and for every $\varepsilon>0$, there exists $T_{0}=T_{0}(t_{0},\varepsilon)$ such that for sufficiently large $n$ and every $3$-graph $\mathcal{H}=(V,E)$ on $n$ vertices, there exists a partition $V=V_{0}\cup V_{1}\cup\cdots\cup V_{L}$ satisfying:
  1. $(i)$
    $L_{0}\leq L\leq T_{0}$;
  2. $(ii)$
    $\left|V_{1}\right|=\cdots=\left|V_{L}\right|$ and $\left|V_{0}\right|\leq\varepsilon n$;
  3. $(iii)$
    for all but at most $\varepsilon\binom{L}{3}$ triples $\left\{i,j,k\right\}\subseteq[L]$, we have that $(V_{i},V_{j},V_{k})_{\mathcal{H}}$ is $\varepsilon$-regular.
The vertex partition in Lemma 3.3 is called an $\varepsilon$-regular partition of $\mathcal{H}$. We call $V_{0}$ the exceptional set, and $V_{1},\ldots,V_{L}$ clusters. For our purpose, we will use the degree form of the weak hypergraph regularity lemma [22], which is proved in the same way as the original regularity lemma of graphs (see proof in [29]).
Lemma 3.4 (Degree form of the weak hypergraph regularity lemma).
For any integer $L_{0}\geq 1$ and every $\varepsilon>0$, there is an integer $n_{0}=n_{0}(\varepsilon,L_{0})$ such that for every $d\in[0,1)$ and for every $3$-graph $\mathcal{H}=(V,E)$ on $n\geq n_{0}$ vertices, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$ and a spanning sub-hypergraph $\mathcal{H^{\prime}}$ of $\mathcal{H}$ such that the following properties hold:
  1. $(i)$
    $L_{0}\leq L\leq n_{0}$ and $\left|V_{0}\right|\leq\varepsilon n$;
  2. $(ii)$
    $\left|V_{1}\right|=\ldots=\left|V_{L}\right|:=m$;
  3. $(iii)$
    $d_{\mathcal{H^{\prime}}}(v)>d_{\mathcal{H}}(v)-(d+\varepsilon)n^{2}$ for each $v\in V$;
  4. $(iv)$
    every edge of $\mathcal{H^{\prime}}$ with more than one vertex in a single cluster $V_{i}$ for some $i\in[L]$ has at least one vertex in $V_{0}$;
  5. $(v)$
    for all triples $\{i,j,k\}\in\binom{L}{3}$, we have that $\left(V_{i},V_{j},V_{k}\right)_{\mathcal{H^{\prime}}}$ is either empty or $(\varepsilon,d)$-regular.
When applying the regularity lemma, a necessary point is to consider the reduced hypergraph. Now, we define the reduced hypergraph whose vertices are the clusters $V_{i}$ and hyperedges correspond to those triples of clusters that induce an $\varepsilon$-regular $3$-graph of sufficient density.
Definition 3.5 (Reduced hypergraph).
Let $\mathcal{H}=(V,E)$ be a $3$-graph. Given parameters $\varepsilon>0$, $d\in[0,1)$ and $L_{0}\geqslant 1$, we define the reduced hypergraph $\mathcal{R}=\mathcal{R}\left(\varepsilon,d,L_{0}\right)$ of $\mathcal{H}$ as follows. Apply the degree form of the weak hypergraph regularity lemma to $\mathcal{H}$, with parameters $\varepsilon,d,L_{0}$ to obtain a spanning sub-hypergraph $\mathcal{H^{\prime}}$ and a partition $V_{0},\ldots,V_{L}$ of $V$. Then $\mathcal{R}$ has vertices $V_{1},\ldots,V_{L}$, and there exists an edge between $V_{i},V_{j},V_{k}$ precisely when $\left(V_{i},V_{j},V_{k}\right)_{\mathcal{H}^{\prime}}$ is $(\varepsilon,d)$-regular.
The reduced hypergraph $\mathcal{R}$ is also a $3$-graph. The following lemma is the case $k=3$ and $\ell=2$ of Lemma 5.5 in [22], which shows that the reduced hypergraph almost inherits the vertex degree of the original hypergraph.
Lemma 3.6 ([22], Lemma 5.5).
Suppose $c>0,L_{0}\geq 1$ and $640<\varepsilon\leq d\leq c^{3}/$ 64. Let $\mathcal{H}$ be an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq cn$. Suppose $\mathcal{R}=\mathcal{R}(\varepsilon,d,L_{0})$ be the reduced hypergraph of $\mathcal{H}$. Then at least $\binom{|V(\mathcal{R})|}{2}-36d^{1/3}|V(\mathcal{R})|^{2}$ of the 2-tuples of vertices of $\mathcal{R}$ have degree at least $(c-4d^{1/3})|V(\mathcal{R})|$.
3.2 Weak 3-graph blow-up lemma
In this subsection, we introduce the transversal blow-up lemma, developed recently by the first author and Staden (Theorem 1.8 in [4]). We will use the $3$-graph version as follows, see Theorem 6.1 in [4]. Before we state the blow-up lemma, some definitions are given.
An $n$-vertex 2-graph $G$ is $\mu$-separable if there is $X\subseteq V(G)$ of size at most $\mu n$ such that the induced subgraph $G[V(G)\setminus X]$ consists of components of size at most $\mu n$. For instance, trees, 2-regular graphs, powers of Hamilton cycles, and graphs with small bandwidths are separable graphs. A 1-expansion $G^{+}$ of a 2-graph $G$ is a $3$-graph obtained by replacing every edge $e=xy$ with a 3-edge $xyt_{e}$, where all the vertices $t_{e}(e\in E(G))$ are distinct and they are disjoint from $V(G)$. We call these added vertices $t_{e}(e\in E(G))$ new vertices of $1$-expansion $G^{+}$. For example, it is easy to see that a loose cycle is a 1-expansion of a cycle.
Theorem 3.7 ([4], Weak $3$-graph blow-up lemma).
Let $0<1/m\ll\varepsilon,\mu,\alpha\ll\nu,d,\delta,1/\Delta,1/r\leq 1$.
  1. $(i)$
    Let $R$ be a 2-graph with vertex set $[r]$.
  2. $(ii)$
    Let $\mathcal{H}$ be a $3$-graph with vertex sets $V_{1},\ldots,V_{r}$ and $V_{ij}$ for $ij\in E(R)$ where $m\leq|V_{i}|\leq m/\delta$ for all $i\in[r]$, and $|V_{ij}|\geq\delta m$ for all $ij\in[r]$. Suppose that $\mathcal{H}[V_{i},V_{j},V_{ij}]$ is weakly $(\varepsilon,d)$-(half)superregular for all $ij\in E(R)$.
  3. $(iii)$
    Let $G$ be a $\mu$-separable 2-graph with $\Delta(G)\leq\Delta$ for which there is a graph homomorphism $\phi:V(G)\rightarrow V(R)$ with $|\phi^{-1}(i)|=|V_{i}|$ for all $i\in[r]$ and $e(G[\phi^{-1}(i),\phi^{-1}(j)])=|V_{ij}|$ for all $ij\in E(R)$.
  4. $(iv)$
    Suppose that for each $i\in[r]$ there is a set $U_{i}\subseteq\phi^{-1}(i)$ with $|U_{i}|\leq\alpha m$ and $T_{x}\subseteq V_{\phi(x)}$ with $|T_{x}|\geq\nu m$ for all $x\in U_{i}$.
Then $\mathcal{H}$ contains a copy of $G^{+}$, where each vertex $x\in V(G)$ is mapped to $V_{\phi(x)}$ and the new vertex $t_{xy}(xy\in E(G))$ is mapped to $V_{\phi(x)\phi(y)}$; and moreover, for every $i\in[r]$ every $x\in U_{i}$ is mapped to $T_{x}$.
Using Lemma 3.7, we prove the following path embedding lemma:
Lemma 3.8 (Path embedding lemma).
Let $0<1/m\ll\varepsilon\ll d\leq 1$ and $m^{\prime}\ll m$. Let $G=G(X_{1},X_{2},X_{3})$ be a $3$-partite $3$-graph and $m\leq\left|X_{i}\right|\leq 2m$ for all $i\in[3]$. If $G\left[X_{1},X_{2},X_{3}\right]$ is weakly $(\varepsilon,d)$-(half)superregular, then for any $V_{i}\subseteq X_{i},V_{j}\subseteq X_{j}$ with $|V_{i}|,|V_{j}|\geq dm$ for any $i,j\in[3]$, $G$ contains a loose path with length $m^{\prime}$ that is no less than $1+0.1|i-j|$ and this path’s endpoints $v_{1},v_{2}$ are located in $V_{i}$ and $V_{j}$.
Proof.
We divide the proof into the following cases according to the length of the path and the values of $i,j$:
Case 1: The path has odd length and $i\neq j$. Let $\{i,j,k\}=\{1,2,3\}.$ Note that there exist a bipartite graph $H^{\prime}$ with vertex set $X_{i}$ and $X_{j}$, such that the number of edges of $H^{\prime}$ equals to $|X_{k}|$ and $H^{\prime}$ contains a path with length $m^{\prime}$ and endpoints $v_{1},v_{2}$, since we can add some vertex-disjoint cycles, $K_{3,3}$-tilings and paths to the $m^{\prime}$-length path such that $|E(H^{\prime})|=|X_{k}|$. Note that $H^{\prime}$ is $\mu$-separable for some $\mu\ll d$, thus using Lemma 3.7, we can embed a $1$-expansion of $H^{\prime}$ into $G$. Moreover, $v_{1},v_{2}$ are embedded into $V_{i}$ and $V_{j}$. Thus, we find a loose path with length $m^{\prime}$, and endpoints $v_{1}\in V_{i}$ and $v_{2}\in V_{j}$.
For other cases, we can generalize the proof.
Case 2: The path has even length and $i\neq j$. For a vertex $v_{1}\in V_{i}$ consider the vertices in $X_{k}$ which together with $v_{1}$ and another vertex in $V_{j}$ forms a hyperedge, which we denoted by $V_{k}$, then embed a path with length $m^{\prime}-1$ and this path’s endpoints $v^{\prime}_{1},v_{2}$ are located in $V_{k}$ and $V_{j}$. This is available since $G$ is weakly $(\varepsilon,d)$-(half)superregular, we have $|V_{k}|\geq d|X_{k}|$, thus there exists a loose path with length $m^{\prime}-1$, and endpoints $v^{\prime}_{1}\in V_{k}$ and $v_{2}\in V_{j}$ follows by a same reason of Case 1, and this path together the edge formed by $v_{1},v^{\prime}_{1}$ and another vertex in $V_{j}$ forms the path we need.
Case 3: The path has even length and $i=j$. Let $\{i,k,l\}=\{1,2,3\}$. For a vertex $v_{1}\in V_{i}$ consider the vertices in $X_{k}$ with $k\neq i$ which together with $v_{1}$ and another vertex in $V_{l}$ forms a hyperedge, which we denoted by $V_{k}$, then embed a path with length $m^{\prime}-1$ and this path’s endpoints $v^{\prime}_{1},v_{2}$ are located in $V_{k}$ and $V_{j}$. This is available since $G$ is weakly $(\varepsilon,d)$-(half)superregular, we have $|V_{k}|\geq d|X_{k}|$, thus there exists a loose path with length $m^{\prime}-1$, and endpoints $v^{\prime}_{1}\in V_{k}$ and $v_{2}\in V_{j}$ follows by a same reason of Case 1, and this path together the edge formed by $v_{1},v^{\prime}_{1}$ and another vertex in $V_{j}$ forms the path we need.
Case 4: The path has odd length and $i=j$. Let $\{i,k,l\}=\{1,2,3\}$. For a vertex $v_{1}\in V_{i}$ consider a 2-length loose path $(v_{1}yz)(zx^{\prime}y^{\prime})$ with $y,y^{\prime}\in X_{k}$, $x\in X_{i}$ and $z\in X_{l}$ , and we denote the vertices in $X_{k}$ with $k\neq i$ that can plays a role as $y^{\prime}$ by $V_{k}$, then embed a path with length $m^{\prime}-2$ and this path’s endpoints $v^{\prime}_{1},v_{2}$ are located in $V_{k}$ and $V_{j}$. This is available since $G$ is weakly $(\varepsilon,d)$-(half)superregular, we have $|V_{k}|\geq d|X_{k}|$, thus there exists a loose path with length $m^{\prime}-2$, and endpoints $v^{\prime}_{1}\in V_{k}$ and $v_{2}\in V_{j}$ follows by a same reason of Case 1, and this path together the $2$-length loose path forms the path we need. ∎
3.3 Basic Probabilistic Estimates
In this subsection, we will introduce a probabilistic tool given by Kühn and Osthus [21]. Given a positive number $\gamma$, set $T$ and sets $A,Q\subseteq T$, we say that $A$ is split $\gamma$-fairly by $Q$ if:
$|\frac{|A\cap Q|}{|Q|}-\frac{|A|}{|T|}|\leqslant\gamma.$
Thus, if $\gamma$ is small and $A$ is split $\gamma$-fairly by $Q$, then the proportion of all those elements of $T$ which lie in $A$ is almost equal to the proportion of all those elements of $Q$ which lie in $A$. The following fact implies that if $Q$ is random, it tends to split large sets $\gamma$-fairly.
Lemma 3.9 ([21], Proposition 4.1).
For each $10<\gamma<$ 1, there exists an integer $q_{0}=q_{0}(\gamma)$ such that the following holds. Given $t\geqslant q\geqslant q_{0}$ and a set $T$ of size $t$, let $Q$ be a subset of $T$ obtained by successively selecting $q$ elements uniformly at random without repetitions. Let $\mathcal{A}$ be a family of at most $q^{10}$ subsets of $T$ such that $|A|\geqslant\gamma t$ for each $A\in\mathcal{A}$. Then, with probability at least $1/2$, every set in $\mathcal{A}$ is split $\gamma$-fairly by $Q$.
3.4 Embedding Cycles in Tripartite Graph
Lemma 3.10.
For any cycle $C$ of size $n$ and any $3$ integers $a,b,c$ with $a+b+c=n$ and $0\leq a\leq b\leq c\leq\lfloor n/2\rfloor$, there exists a proper $(a,b,c)$-coloring of $C$, i.e., $C$ has a proper vertex-coloring such that the three monochromatic parts have sizes $a,b$ and $c$.
Proof.
Without loss of generality, suppose that the vertex set of $C$ is $\mathbb{Z}_{n}=\mathbb{Z}/n\mathbb{Z}$ and the edge set of $C$ is $\{(i,i+1)\mid i\in\mathbb{Z}_{n}\}$. We partition the vertex $V(C)$ into two parts $A$ and $B$ where $A=\{i\in\mathbb{Z}_{n}\mid 1\leq i\leq n,i\ \text{is odd}\}$ and $B=\{i\in\mathbb{Z}_{n}\mid 1\leq i\leq n,i\ \text{is even}\}$. We will color $C$ by using three different colors red, blue, and green as follows. First, note that $|A|=\lceil n/2\rceil$ and $c\leq\lceil n/2\rceil$, thus we have $\{1,3,\ldots,2c-1\}\subseteq A$ and we color these vertices by red. Let $A^{\prime}=A\backslash\{1,3,\ldots,2c-1\}$. If $|A^{\prime}|\geq b$, then this implies that $a\geq|B|=\lfloor n/2\rfloor$ and thus $a=\lfloor n/2\rfloor$ and $|A^{\prime}|=b$ by our assumption. In such case, we color all the vertices in $A^{\prime}$ by blue and color all the vertices in $B$ by green. It is easy to check that no two monochromatic vertices in $C$ are adjacent.
Therefore, we suppose that $|A^{\prime}|<b$. We color all the vertices in $A^{\prime}$ by blue and color the set $\{2,4,\ldots,2(b-|A^{\prime}|)\}$ by blue. We then color all the remaining vertices in $B$ by green. We claim that under such a rule, there exists no adjacent monochromatic pair in $C$. Two adjacent monochromatic vertex pairs could only exist in the blue vertex set, and this will imply that $2(b-|A^{\prime}|)\geq 2c$. Recall that $|A^{\prime}|=\lfloor n/2\rfloor-c$, combining with the previous inequality, we obtain $b\geq\lfloor n/2\rfloor$ and thus $b=c=n/2$ and $n$ is even by our assumption. But if so, $A^{\prime}$ is empty and the whole vertex set $A$ is red. Besides, by our coloring rule, the whole vertex set $B$ is green. Hence there is no adjacent monochromatic pair in $C$ and thus completes the proof of the lemma. ∎
Lemma 3.11.
Let $F$ be a graph consisting of $r$ vertex-disjoint cycles of order $n_{1},n_{2},\ldots,n_{r}$ satisfying $n_{1}+n_{2}+\cdots+n_{r}=n$ and the number of odd cycles is $k$, and $H=(X,Y,Z)$ be a complete tripartite graph on $n$ vertices with minimum degree at least $\sum_{i=1}^{r}\lceil n_{i}/2\rceil=n/2+k/2$, then $H$ contains $F$ as a subgraph.
Proof.
Without loss of generality, we suppose that $|X|\leq|Y|\leq|Z|$. We prove the lemma by induction on $r$. When $r=1$, suppose first that $H$ is a cycle of even order. By the minimum degree condition, we get $|X|+|Y|\geq n/2$ and thus $|Z|\leq n/2$. Similarly, we get $|X|,|Y|\leq n/2$. Thus we can embed $F$ into $H$ by Lemma 3.10 since $F$ has a proper $(|X|,|Y|,|Z|)$-coloring. Therefore we may suppose that $|F|$ is odd, the proof is similar to the even case, so we omit it.
Suppose that the conclusion is true for $r-1(r\geq 2)$, we prove that it is true for $r$. Suppose $H$ is the disjoint union of cycles $C_{1},\ldots,C_{r}$. We delete exactly $k-1$ vertices from each set $X,Y,Z$ and denote the remaining sets by $X^{\prime},Y^{\prime},Z^{\prime}$. Let $H^{\prime}$ be its induced tripartite graph in $H$.
Claim 3.12.
$|X^{\prime}|+|Y^{\prime}|\geq\lceil n_{1}/2\rceil$.
Proof.
Suppose not, we get $|X^{\prime}|+|Y^{\prime}|<\lceil n_{1}/2\rceil$ and hence $|X|+|Y|<2k-2+\lceil n_{1}/2\rceil$. By the minimum degree condition, we get $|X|+|Y|\geq(n+k)/2$ and hence $(n+k)/2<2k-2+\lfloor n_{1}/2\rfloor$, which implies $n_{2}+\ldots+n_{r}<3k-4$ when $n_{1}$ is even and $n_{2}+\ldots+n_{r}<3k-3$ when $n_{1}$ is odd. This is a contradiction since $n_{2}+\ldots+n_{r}\geq 3k$ when $n_{1}$ is even and $n_{2}+\ldots+n_{r}\geq 3k-3$ when $n_{1}$ is odd. ∎
We divide the remaining proof into the following two cases:
Case 1: $|Y^{\prime}|\geq\lfloor n_{1}/2\rfloor$; By Lemma 3.10, $C_{1}$ has a $(0,\lfloor n_{1}/2\rfloor,\lfloor n_{1}/2\rfloor)$-coloring when $n$ is even and a $(1,\lfloor n_{1}/2\rfloor,\lfloor n_{1}/2\rfloor)$-coloring when $n$ is odd.
Denote the three parts in the ordering by $D_{1},D_{2},D_{3}$ and hence we can trivially embed $C_{1}$ into $H^{\prime}$ by embedding $D_{1},D_{2},D_{3}$ into $X^{\prime},Y^{\prime},Z^{\prime}$ respectively.
Now we delete $C_{1}$ from the graph $H$, denote the resulting graph by $H_{1}$ and its three parts by $X_{1},Y_{1},Z_{1}$. We claim that $\delta(H_{1})\geq(|H_{1}|+k)/2$ when $n$ is even and $\delta(H_{1})\geq(|H_{1}|+k-1)/2$ when $n$ is odd. This is equivalent to that $|S_{1}|\leq(|S_{1}|-k)/2$ when $n$ is even and $|S_{1}|\leq(|S_{1}|-k+1)/2$ when $n$ is odd for $S\in\{X,Y,Z\}$, which follows from $|S_{1}|\leq(n-k)/2-\lfloor n_{1}/2\rfloor$ for $S\in\{X_{1},Y_{1},Z_{1}\}$.
Hence by induction on $H_{1}$, we can embed cycles $C_{2},\ldots,C_{r}$ into $H_{1}$. This combined with $C_{1}$ will complete the proof.
Case 2: $|Y^{\prime}|<\lfloor n_{1}/2\rfloor$; It follows that $|X^{\prime}|+|Y^{\prime}|<n_{1}$ and hence $n_{1}-|X^{\prime}|-|Y^{\prime}|>0$. By Claim 3.12, we get $n_{1}-|X^{\prime}|-|Y^{\prime}|\leq\lfloor n_{1}/2\rfloor$. Therefore, by Lemma 3.10, the cycle $C_{1}$ has a proper $(|X^{\prime}|,|Y^{\prime}|,n_{1}-|X^{\prime}|-|Y^{\prime}|)$-coloring. Recall that we have $|Z^{\prime}|\geq n_{1}-|X^{\prime}|-|Y^{\prime}|$ since $n_{2}+\ldots+n_{r}\geq 3k-3$. Denote the three parts in the ordering by $D_{1},D_{2},D_{3}$ and hence we can trivially embed $C_{1}$ into $H^{\prime}$ by embedding $D_{1},D_{2},D_{3}$ into $X^{\prime},Y^{\prime},Z^{\prime}$ respectively.
Delete $C_{1}$ from the graph $H$, denote the resulting graph by $H_{1}$ and its three parts by $X_{1},Y_{1},Z_{1}$. Similar to Case 1, we get $\delta(H_{1})\geq(|H_{1}|+k)/2$ when $n$ is even and $\delta(H_{1})\geq(|H_{1}|+k-1)/2$ when $n$ is odd. This follows from $|S_{1}|\leq(n-k)/2-\lfloor n_{1}/2\rfloor$ for $S\in\{X_{1},Y_{1},Z_{1}\}$.
Therefore, by induction on $H_{1}$, we can embed cycles $C_{2},\ldots,C_{r}$ into $H_{1}$ and thus complete the proof.
Note that $|G_{1}|\geq 3k$ since each odd cycle has length of at least $3$.
3.5 Definition of some parameters
Definition 3.13.
Let $n>k>0$ be positive integers and $10<\eta<$ 1 be a positive constant. Let $a,b$ be positive integers. We say $a,b$ are $(n,k,\eta)$-good if the following holds:
  • (i)
    $\frac{2n-4k-0.08\eta n}{n+2k+0.04\eta n}\leq\frac{b}{a}\leq\frac{2n-4k}{n+2k}-\eta^{2}$;
  • (ii)
    $a,b\leq\lceil\frac{200}{\eta}\rceil$.
Fact 3.14.
Let $0<\frac{1}{n}<\frac{1}{k}$ and $1/n\ll\eta\ll 1$. There exist positive integers $a=a(\eta),b$ such that $a,b$ are $(n,k,\eta)$-good.
Proof.
First note that the following holds:
$\frac{2n-4k}{n+2k}-\eta^{2}-\frac{2n-4k-0.08\eta n}{n+2k+0.04\eta n}$ $\geq\frac{0.08\eta n}{n+2k+0.04\eta n}-\eta^{2}$
$\geq\frac{0.08\eta}{3+0.04\eta}-\eta^{2}\geq 0.02\eta.$
Thus, when $\eta$ is fixed, choose $a=\lceil\frac{100}{\eta}\rceil$. Since $\big{|}[\frac{2n-4k-0.08\eta n}{n+2k+0.04\eta n},\frac{2n-4k}{n+2k}-\eta^{2}]\big{|}\geq 0.02\eta$ there exists a positive integer $b>0$ such that $b/a\in[\frac{2n-4k-0.08\eta n}{n+2k+0.04\eta n},\frac{2n-4k}{n+2k}-\eta^{2}].$ Moreover $b\leq a\cdot(\frac{2n-4k}{n+2k}-\eta^{2})\leq\lceil\frac{200}{\eta}\rceil$. ∎
4 Proof of Theorem 1.4
In this section, we give a proof of Theorem 1.4. In practice, we reduce the problem to the case that each loose cycle in $\mathcal{C}$ has at most $\log n$ vertices, i.e., Theorem 4.1. Throughout this paper, we use the label log to denote the base 2 logarithm. The case for some loose cycles in $\mathcal{C}$ with larger than $\log n$ vertices is proved in Appendix A.1. This approach draws inspiration from the work of Khan [16]. Thus, the goal of this section is to prove that:
Theorem 4.1.
For every $\eta>0$, there exists an integer $n_{0}=n_{0}(\eta)$ such that the following holds for any $n\geq n_{0}$. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$, where $\sum_{i=1}^{r}n_{i}=n$, each loose cycle has a length of at least $3$, and $n_{i}\leq\log n$ for each $i\in[r]$. Let $k$ denote the number of loose cycles with odd lengths. If $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq(n+2k)/4+\eta n$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph.
We prove these two special cases before the main proof:
Lemma 4.2.
For every $\eta>0$, there exists an integer $n_{0}=n_{0}(\eta)$ such that the following holds for any $n\geq n_{0}$. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$, where $\sum_{i=1}^{r}n_{i}=n$, each loose cycle has a length of at least $3$, and $n_{i}\leq\log n$ for each $i\in[r]$. Let $k$ denote the number of loose cycles with odd lengths. Additionally, there exist an integer $s$ such that the number of $C_{s}$ in the list is at least $(1-0.5\eta)\frac{n}{s}$. If $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq(n+2k)/4+\eta n$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph.
Proof.
By assumption, the number of $C_{s}$ in the list is at least $(1-0.5\eta)\frac{n}{s}$, thus the number of vertices that is not in $C_{s}$ of the list is no more than $n-s\cdot(1-0.5\eta)\frac{n}{s}=0.5\eta n.$ Therefore, by Lemma A.3, we can embed every cycle that is not a $C_{s}$ greedily in $\mathcal{H}$. Delete these embedded cycles from $\mathcal{H}$ and let $\mathcal{H}_{1}$ be the resulting $3$-graph. By Theorem 1.3, the remainder of $\mathcal{H}$ has a $C_{s}$-factor, which completes the proof. ∎
Thus, if $k\geq(1-0.2\eta)\frac{n}{6}$, denote $c_{6}$ as the number of $C_{6}$. We have $6c_{6}+10(k-c_{6})\leq n$, thus $c_{6}\geq(1-0.5\eta)\frac{n}{6}$ and by Lemma 4.2 we have proved the theorem.
Lemma 4.3.
For every $\eta>0$, there exists an integer $n_{0}=n_{0}(\eta)$ such that the following holds for any $n\geq n_{0}$. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$, where $\sum_{i=1}^{r}n_{i}=n$, each loose cycle has a length of at least $3$, and $n_{i}\leq\log n$ for each $i\in[r]$. Let $c_{8}$ denote the number of $C_{8}$ and $c_{12}$ denote the number of $C_{12}$, and $16c_{8}+24c_{12}\geq(1-\eta^{2})n$. If $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq n/4+\eta n$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph.
Proof.
If $c_{8}$ or $c_{12}$ is no more than $0.01\eta$, then this could be proved by a similar argument as in Lemma 4.2 using Theorem 1.3.
Otherwise, by Lemma A.3, we can embed every cycle that is not a $C_{8}$ or $C_{12}$ greedily in $\mathcal{H}$. Delete these embedded cycles from $\mathcal{H}$ and let $\mathcal{H}_{1}$ be the resulting $3$-graph. We randomly partition $V(\mathcal{H}_{1})$ into $V_{1}$ and $V_{2}$ with $|V_{1}|=16c_{8}$ and $|V_{2}|=24c_{12}$. By a similar argument in the Appendix, there is a partition $V=V_{1}\cup V_{2}$ such that $\delta_{2}(\mathcal{H}[V_{1}])\geq(1/4+\eta/3)|V_{1}|$ and $\delta_{2}(\mathcal{H}[V_{2}])\geq(1/4+\eta/3)|V_{2}|$. Then, $C_{8}$ can be embedded into $\mathcal{H}[V_{1}]$ and $C_{12}$ can be embedded into $\mathcal{H}[V_{2}]$ by Theorem 1.3. ∎
Throughout the rest of the paper, we fix positive constants satisfying the following hierarchy:
$0<1/n\ll 1/m\ll 1/L_{0}\ll\varepsilon\ll d\ll\psi\ll\eta\ll 1$
Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$ where $\sum_{i=1}^{r}n_{i}=n$ and $6\leq n_{i}\leq\log n$ for each $i\in[r]$. Let $k$ be the number of loose cycles with odd lengths in $\mathcal{C}$, and let $\mathcal{H}$ be an $n$-vertex $3$-graph with minimum codegree at least $(n+2k)/4+\eta n$.
4.1 Step 1
We first apply the weak regularity lemma (Lemma 3.4) with parameters $n,m,\varepsilon,L_{0}$. This lemma implies that there exists a partition of $V(\mathcal{H})$ into $V_{0},V_{1},\ldots,V_{L}$, where $\left|V_{0}\right|\leq\varepsilon n$ with $\left|V_{1}\right|=\ldots=\left|V_{L}\right|=:m$ and a spanning sub-hypergraph $\mathcal{H}^{\prime}$ of $\mathcal{H}$ such that the following holds:
  1. $(i)$
    $L_{0}\leq L\leq n_{0}$ and $\left|V_{0}\right|\leq\varepsilon n$;
  2. $(ii)$
    $\left|V_{1}\right|=\ldots=\left|V_{L}\right|:=m$;
  3. $(iii)$
    $d_{\mathcal{H^{\prime}}}(v)>d_{\mathcal{H}}(v)-(d+\varepsilon)n^{2}$ for each $v\in V$;
  4. $(iv)$
    every edge of $\mathcal{H^{\prime}}$ with more than one vertex in a single cluster $V_{i}$ for some $i\in[L]$ has at least one vertex in $V_{0}$;
  5. $(v)$
    for all triples $\{i,j,k\}\in\binom{L}{3}$, we have that $\left(V_{i},V_{j},V_{k}\right)_{\mathcal{H^{\prime}}}$ is either empty or $(\varepsilon,d)$-regular.
The reduced hypergraph $\mathcal{R}$ has vertices $V_{1},\ldots,V_{L}$. In this reduced hypergraph, an edge exists between $V_{i_{1}},V_{i_{2}},V_{i_{3}}$ precisely when $\left(V_{i_{1}},V_{i_{2}},V_{i_{3}}\right)_{\mathcal{H}^{\prime}}$ is $(\varepsilon,d)$-regular. By Lemma 3.6, all but at most $36d^{1/3}L^{2}$ of the $2$-tuples of vertices in $\mathcal{R}$ have degree at least $\left(\frac{n+2k}{4n}+\eta-4d^{1/3}\right)L$.
Let $q>p$ be positive integers. Define the $3$-graph $\mathcal{A}_{p,q}^{3}$ as follows. The vertex set $V\left(\mathcal{A}_{p,q}^{3}\right)$ is the union of disjoint sets $A_{1},\ldots,A_{q-p}$ and $B$, where $\left|A_{j}\right|=2$ for each $j\in[q-p]$ and $|B|=2p$. Then any $3$-tuple of the form $\{x\}\cup A_{j}$ with $x\in B$ is an edge of $\mathcal{A}_{p,q}^{3}$ and $j\in\left[q-p\right]$. In particular, we have $\left|V\left(\mathcal{A}_{p,q}^{3}\right)\right|=2q$. Mycroft [24] gave the following result. In fact, he provided a more general result for all uniform hypergraphs.
Lemma 4.4 ([24]).
Suppose that $31/m\ll\theta\ll\psi\ll 1/q,1/p,1/$ 3 and that $G$ is a $3$-graph on vertex set $[m]$ such that $\operatorname{deg}_{G}(S)>\left(\frac{p}{q}+\theta\right)m$ for all but at most $\theta m^{2}$ sets $S\in\binom{m}{2}$. Then $G$ admits an $\mathcal{A}_{p,q}^{3}$-tiling $\mathcal{F}$ such that $|V(\mathcal{F})|\geq(1-\psi)m$.
Let $a,b$ be integers such that $a,b$ are $(n,k,\eta)$-good. Thus, $\frac{a}{2a+b}=\frac{1}{2+b/a}\leq\frac{n+2k}{n}+0.01\eta n$. Hence for all but at most $36d^{1/3}L^{2}$ of the $2$-tuples of vertices in $\mathcal{R}$ have degree at least
$(\frac{n+2k}{4n}+\eta-4d^{1/3})L\geq(\frac{a}{2a+b}+0.5\eta-4d^{1/3})L>(\frac{a}{2a+b}+36d^{1/3})L.$
Choose $p=a,q=2a+b$ and apply Lemma 4.4 to the reduced hypergraph $\mathcal{R}$ with $\theta=36d^{1/3}\ll\eta$ and thus $1/L\ll\theta\ll\psi\ll\eta,1/a,1/b$. Thus, we have the following claim:
Claim 4.5.
There exists an $\mathcal{A}_{a,2a+b}^{3}$-tiling $\mathcal{M}$ in $\mathcal{R}$ such that all but at most $\psi L$ vertices in $\mathcal{R}$ are covered by $\mathcal{M}$.
By putting at most $\psi Lm\leq\psi n$ vertices into $V_{0}$ from the clusters that are not covered by $\mathcal{M}$, we get an $\mathcal{A}_{a,2a+b}^{3}$-factor in $\mathcal{R}$. Here, we still denote the $\mathcal{A}_{a,2a+b}^{3}$-factor by $\mathcal{M}$. Now, we have $|V_{0}|\leq(\varepsilon+\psi)n$.
For each element $\mathcal{A}_{a,2a+b}^{3}\in\mathcal{M}$, denote the clusters in $\mathcal{A}_{a,2a+b}^{3}$ by $A_{i,1},A_{i,2}$ for every $i\in\left[a+b\right]$ and $B_{j}$ for $j\in[2a]$. Let $B$ be the union of $B_{j}$ for $j\in\left[2a\right]$. For each hyperedge $A_{i,1}A_{i,2}B$ of $\mathcal{A}_{a,2a+b}^{3}$, the $3$-partite $3$-graph $(A_{i,1},A_{i,2},B)_{\mathcal{H}^{\prime}}$ is $(2a\varepsilon,d/2a)$-half-regular, since for any $A_{1}\subseteq A_{i,1},A_{2}\subseteq A_{i,2}$ and $B_{0}\subseteq B$ with $|A_{1}|\geq 2a\varepsilon|A_{i,1}|,|A_{2}|\geq 2a\varepsilon|A_{i,2}|$ and $|B_{0}|\geq 2a\varepsilon|B|$ we have:
$d(A_{1},A_{2},B_{0})=\sum_{j\in[2a]}e(A_{1},A_{2},B_{0}\cap B_{j})/|A_{1}||A_{2}||B_{0}|\geq\frac{d\cdot|A_{1}|\cdot|A_{2}|}{|A_{1}|\cdot|A_{2}|\cdot 2a}=d/2a,$
where the inequality holds because when $|B_{0}\cap B_{j}|\geq\varepsilon|B_{j}|$ we have $e(A_{1},A_{2},B_{0}\cap B_{j})\geq d\cdot|A_{1}|\cdot|A_{2}|\cdot|B_{0}\cap B_{j}|$; otherwise, we trivially have $e(A_{1},A_{2},B_{0}\cap B_{j})\geq 0$. When $|B_{0}|\geq 2a\varepsilon|B|$ there exists $j^{\prime}\in[2a]$ such that $|B_{0}\cap B_{j^{\prime}}|\geq|B_{0}|/2a\geq\varepsilon|B_{j^{\prime}}|$, thus $\sum_{j\in[2a]}e(A_{1},A_{2},B_{0}\cap B_{j})\geq e(A_{1},A_{2},B_{0}\cap B_{j^{\prime}})\geq d\cdot|A_{1}|\cdot|A_{2}|\cdot|B_{0}\cap B_{j^{\prime}}|$ and $|B_{0}\cap B_{j^{\prime}}|/|B_{0}|\geq 1/2a$.
By moving at most $(a+b)$ vertices into $V_{0}$ from each cluster if required, we can ensure that the number of vertices in each cluster is divisible by $(a+b)$. At this point, we define the size of each cluster as $m_{1}$. By Lemma 3.1, the $3$-partite $3$-graph $(A_{i,1},A_{i,2},B)_{\mathcal{H^{\prime}}}$ is $(4a\varepsilon,d/4a)$-half-regular by this step. Then we partition the clusters to get many disjoint triples that cover almost all vertices of $\mathcal{H}$. The partition goes as follows:
We partition $B$ into $2(a+b)$ sets with equal size uniformly at random. We label these equal parts as $Y_{i,1},Y_{i,2}$ for every $i\in\left[a+b\right]$. Meanwhile, for each $i\in[a+b]$, we partition every $A_{i,1}$ into two subsets $X_{i,1}$ and $Z_{i,2}$ and partition every $A_{i,2}$ into two subsets $X_{i,2}$ and $Z_{i,1}$ uniformly at random such that $|X_{i,1}|:|Z_{i,2}|=|X_{i,2}|:|Z_{i,1}|=a:b$. For each $\mathcal{A}_{a,2a+b}^{3}\in\mathcal{M}$, after the random partitioning, we obtain $2(a+b)$ triples $(X_{i,j},Y_{i,j},Z_{i,j})$ for $i\in[a+b]$ and $j\in\{1,2\}$ where $|X_{i,j}|=|Y_{i,j}|=\frac{am_{1}}{a+b}$ and $|Z_{i,j}|=\frac{bm_{1}}{a+b}$. Let $N:=2(a+b)|\mathcal{M}|$ denote the number of such triples, and take relabel them by $(X_{i},Y_{i},Z_{i})$ for each $i\in[N]$.
By Lemma 3.1, the $3$-partite $3$-graph $(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ is $(8a\varepsilon,d/8a)$-half-regular for every $i\in[N]$. Moreover, we assert that the following claim holds, and its statement and proof are similar to Proposition 7.1 in [21]:
Claim 4.6.
Let $\eta_{1}=\eta-4d^{1/3}$ and $\eta^{\prime}_{1}=0.99\eta_{1}$. For every pair of vertices $u,v\in V(\mathcal{H})$, there are at least $\eta^{\prime}_{1}\frac{N}{2(2a+b)}$ indices $i$ for which
$|N_{\mathcal{H}}(u,v)\cap Z_{i}|\geq\eta^{\prime}_{1}\frac{|Z_{i}|}{2(2a+b)}.$
Proof.
For any cluster $V_{i}$ ($i\in[L]$), we say it is $(u,v)$-good if $|N_{\mathcal{H}}(u,v)\cap V_{i}|\geq\eta_{1}\frac{m}{2((2a+b)-a)}.$ Let $N^{\prime}_{uv}$ denote the number of elements in the $\mathcal{A}_{a,(2a+b)}$-tiling $\mathcal{M}$ which contain at least $12a+$ 1 $(u,v)$-good clusters. Then
$\frac{n+2k}{4}+\eta_{1}n\leq|N_{\mathcal{H}}(v,w)|\leq 2(2a+b)mN^{\prime}_{uv}+|\mathcal{M}|\left(2am+2((2a+b)-a)\frac{\eta_{1}m}{2(2a+b)-2a}\right).$
Since $n\geq 2(2a+b)|\mathcal{M}|m$, this implies that $N^{\prime}\geq\frac{(2(2a+b)-1)\eta_{1}|\mathcal{M}|}{2(2a+b)}$. Consider any $\mathcal{A}_{a,(2a+b)}\in\mathcal{M}$ which contains at least $12a+$ 1 (original) clusters $V_{t}$ such that $|V_{t}\cap N_{\mathcal{H}}(v,w)|\geq\eta_{1}\frac{m}{2(2a+b)-2a}$. Thus, for some $A_{i}$ in this $\mathcal{A}_{a,(2a+b)}$, $|A_{i}\cap N_{\mathcal{H}}(v,w)|\geq\eta_{1}\frac{m}{2(2a+b)-2a}-(a+b)$. However, $A_{i}$ contains some set of the form $Z_{i}$. Lemma 3.9 implies that we may assume that for the pair $v,w$, the intersection of $N_{\mathcal{H}}(v,w)$ with this $Z_{i}$ has size at least $\eta_{1}\frac{|Z_{i}|}{2(2a+b)}-(a+b)>\eta^{\prime}_{1}\frac{|Z_{i}|}{2(2a+b)}$. Thus, we have found at least $N^{\prime}_{uv}\geq\eta^{\prime}_{1}\frac{N}{2(2a+b)}$ sets $Z_{i}$, each containing at least $\eta^{\prime}_{1}\frac{|Z_{i}|}{2(2a+b)}$ neighbors of the pair $v,w$, as required. ∎
By Lemma 3.2, after removing at most $\varepsilon m^{\prime}L\leq\varepsilon n$ vertices into $V_{0}$, we obtain that the $3$-partite $3$-graph $(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ is $(16a\varepsilon,d/16a)$-half-superregular for every $i\in[N]$. Then for each $i\in[N]$ remove $\sqrt{m}$ vertices from $Z_{i}$ to $V_{0}$. By Lemma 3.1 the $3$-partite $3$-graph $(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ is $(32a\varepsilon,d/32a)$-half-superregular for every $i\in[N]$. We set $\varepsilon_{1}=32a\varepsilon$ and $d_{1}=d/32a$.
For each $i\in[N]$, we have $|Z_{i}|+\sqrt{m}=\frac{b}{a}|X_{i}|=\frac{b}{a}|Y_{i}|$ and for every $i\in[N]$ let $T_{i}:T_{i}(X_{i},Y_{i},Z_{i})=(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ denote these triples. Since if $k\geq(1-0.2\eta)\frac{n}{6}$ by Lemma 4.2 it is proved, without loss of generality we have $b/a\geq\frac{2n-4k-0.04\eta n}{n+2k+0.02\eta n}>1+\eta^{2}$, thus
$|Z_{i}|-|X_{i}|\geq(1-\varepsilon)m_{1}(\frac{b}{a+b}-\frac{a}{a+b})-\sqrt{m}\geq(1-\varepsilon)m_{1}(\eta^{2}\frac{a}{a+b})-\sqrt{m}\geq(1+\eta^{3})m_{1}.$
Let $\eta^{\prime}=\eta^{\prime}_{1}-2\varepsilon$. Thus, for every pair of vertices $u,v\in V(\mathcal{H})$, there are at least $\eta^{\prime}\frac{N}{2(2a+b)}$ indices $i$ for which
$|N_{\mathcal{H}}(u,v)\cap Z_{i}|\geq\eta^{\prime}\frac{|Z_{i}|}{2(2a+b)}.$
The number of vertices in $V_{0}$ is at most
$|V_{0}|\leq(2\varepsilon+\psi)n+(a+b)L+N\sqrt{m}.$
Let $\xi=\big{(}(2\varepsilon+\psi)n+(a+b)L+N\sqrt{m}\big{)}/n$, we have $|V_{0}|\leq\xi n$ and $\xi\ll\eta^{\prime},\eta.$ Now, almost all vertices, except for the vertices in $V_{0}$, are covered by triples $T_{1},T_{2},\ldots,T_{N}$.
4.2 Step 2
Now, we show the detailed embedding program. To begin with, we embed some loose cycles to use up all the vertices in $V_{0}$. Let $x$ be the smallest integer that satisfies the following inequality:
$\sum_{i=1}^{x}\frac{n_{i}}{2}\geq|V_{0}|.$
Remove exactly $\sum_{i=1}^{x}\frac{n_{i}}{2}-|V_{0}|$ vertices from some $Z_{i}$ into $V_{0}$ such that
$|V_{0}|=\sum_{i=1}^{x}\frac{|V(C_{n_{i}})|}{2}.$
Let $v_{1},v_{2},\ldots,v_{t}$ be an enumeration of the vertices in $V_{0}$. We will embed the loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{x}}$ to use up all the vertices in $V_{0}$.
To embed the cycle $C_{n_{1}}$ to cover vertices $v_{1},v_{2},\ldots,v_{n_{1}/2}$, we claim that for each pair $v_{i}v_{i+1}$ (and $v_{1}v_{n_{1}/2}$) for $i\in[n_{1}/2-1]$, we can find distinct vertices $w$ such that $v_{i}wv_{i+1}$ (and $v_{1}wv_{n_{1}/2}$) is a hyperedge of $\mathcal{H}$, and such that $w$ is contained in one of the sets $Z_{i}$, and moreover, this hold for every $C_{n_{i}}(i\in[x])$. This can be obtained by Claim 4.6. Since $|V(T_{i})|\leq 3m$ for each $i\in[N]$, the total number of vertices $w$ that we will use is at most
$|V_{0}|\leq\xi n\leq 4\xi Nm.$
On the other hand,
$|V_{0}|\leq\xi n\leq 4\xi Nm\leq\left(\sqrt{\xi}m\right)\left(\frac{\eta^{\prime}N}{2(2a+b)}\right),$
which follows from $\xi\ll\eta^{\prime}$ and $a,b\leq 200/\eta+1\leq 250/\eta^{\prime}$. Thus, by Lemma 4.6, we can choose the vertices $w$ greedily such that none of the $Z_{i}$ has been consumed by more than $\sqrt{\xi}m$ vertices. To ensure this, we randomly select a subset $Z^{\prime}_{i}\subseteq Z_{i}$ of size $\eta^{4}|Z_{i}|$ for each $Z_{i}$ and only use the vertices in $Z^{\prime}_{i}$ to play the role of $w$. We remove all the vertices $w$ from the sets $Z_{i}$ and still denote the resulting sets by $Z_{i}$. We claim that having finished this $(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ is sufficiently half-superregular for each $i\in[N]$ by removing at most $\sqrt{\xi}m$ vertices from each $Z_{i}$:
Claim 4.7.
Let $\varepsilon^{*}:=1.01\varepsilon_{1}$ and $d^{*}:=d_{1}^{2}/16$. The resulting $3$-partite $3$-graph $(X_{i},Y_{i},Z_{i})_{\mathcal{H}^{\prime}}$ is $(\varepsilon^{*},d^{*})$-half-superregular for $i\in[N]$.
Proof.
For each $Z_{i}$, we randomly select a subset denoted as $Z^{\prime}_{i}\subseteq Z_{i}$ with a size of $\eta^{4}|Z_{i}|$. By Lemma 3.1, with high probability, the triple $(X_{i},Y_{i},Z_{i}\setminus Z^{\prime}_{i})_{\mathcal{H^{\prime}}}$ is $(1.01\varepsilon_{1},d_{1}^{2}/16)$-superregular. Thus by union bound, there exist a such selection such that $(X_{i},Y_{i},Z_{i}\setminus Z^{\prime}_{i})_{\mathcal{H^{\prime}}}$ is $(1.01\varepsilon_{1},d_{1}^{2}/16)$-superregular for every $i\in[N]$. According to Lemma 3.9, we claim that for every pair of vertices $v$ and $w$ in $V(H)$, there exist at least $\eta^{\prime}N/2(2a+b)$ indices $i$ such that
$|N_{H}(v,w)\cap Z^{\prime}_{i}|\geq\frac{\eta^{\prime}|Z^{\prime}_{i}|}{3(2a+b)}\gg\sqrt{\xi}m.$
Thus, for every $i\in[N]$ we can greedily select no more than $\sqrt{\xi}m$ vertices from $Z^{\prime}_{i}$ to form cycles together with the vertices in $V_{0}$, and then remove these vertices from $Z_{i}^{\prime}$. Then put every vertex remained in $Z^{\prime}_{i}$ to $Z_{i}$, and still denote the triple as $(X_{i},Y_{i},Z_{i})$.
Since for every subsets $X_{i}^{\prime\prime}\subseteq X_{i},Y_{i}^{\prime\prime}\subseteq Y_{i},Z_{i}^{\prime\prime}\subseteq Z_{i}$ with $|X_{i}^{\prime\prime}|\geq 1.01\varepsilon_{1}|X_{i}|,|Y_{i}^{\prime\prime}|\geq 1.01\varepsilon_{1}|Y_{i}|,|Z_{i}^{\prime\prime}|\geq 1.01\varepsilon_{1}|Z_{i}|\geq\varepsilon_{1}\frac{bm_{1}}{a+b}$, $d(X_{i}^{\prime\prime},Y_{i}^{\prime\prime},Z_{i}^{\prime\prime})\geq d_{1}\geq d_{1}^{2}/16$, $(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ is $(1.01\varepsilon_{1},d_{1}^{2}/16)$-half-regular for every $i\in[N]$. Moreover, for any vertex in $Z_{i}$, its degree is at least $d_{1}$, and for any vertex from $X_{i}$ or $Y_{i}$, its degree is at least $d_{1}^{2}/16$ by Lemma 3.1. Thus the claim holds. ∎
We denote these triples as $T_{i}=(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$. It satisfies the following for every $i\in[N]$:
$0<\frac{1}{n}<\frac{1}{|T_{i}|}\ll\varepsilon^{*}\ll d^{*}\ll\xi\ll\eta\ll 1$
For every $X_{i}$ and $Y_{j}$, we are going to find some vertices together with one vertex in $X_{i}$ and another in $Y_{j}$ to form a hyperedge. The following Claim shows that there is some connectivity between any two triples. In fact, $\sqrt{n}$ can be improved to linear $n$, but $\log n$ is enough for our needs here.
Claim 4.8.
For any $X_{i}$ and $Y_{j}$, there exists a triple $T_{h}$ and $X^{0}_{i}\subseteq X_{i}$ with $|X^{0}_{i}|\geq 0.001|X_{i}|$ such that the following holds:
For any $x_{i}\in X^{0}_{i}$, there exists at least $0.001|Y_{j}|$ vertices $y_{j}\in Y_{j}$ such that for every such $y_{j}$, there exists at least $\sqrt{n}$ numbers of $t_{j}$ in $T_{h}$ with $(x_{i}t_{j}y_{j})\in E(\mathcal{H})$.
Proof.
Let $\widetilde{T}=\cup T_{i}$ and $X_{i}Y_{j}=\{xy:\forall x\in X_{i},y\in Y_{j}\}$. Consider a bipartite graph $B$ with vertex set $V\left(X_{i}Y_{j},V(\widetilde{T})\right)$. An edge exists between a vertex pair $xy$ and a vertex $t\in\widetilde{T}$ in the bipartite graph if and only if $(xyt)\in E(\mathcal{H})$. According to the codegree condition, we know that in the bipartite graph $B$, every vertex in the set $X_{i}Y_{j}$ has degree $d\geq\left(\frac{1}{4}+\eta-2\xi\right)n-\log n\geq\frac{1}{4}n\geq\frac{1}{4}\left|\widetilde{T}\right|$. Now, we claim that the number of vertices in $\widetilde{T}$ with degree no less than $\frac{1}{10}\left|X_{i}Y_{j}\right|$ is at least $\frac{n}{7}$. To prove this, assume the contrary. By double counting, we have:
$\frac{n}{4}\left|X_{i}Y_{j}\right|$ $\leq E(B)$
$\leq\frac{1}{10}\left|X_{i}Y_{j}\right|\cdot n+\left|X_{i}Y_{j}\right|\cdot\frac{1}{7}n$
$<\frac{n}{4}\left|X_{i}Y_{j}\right|.$
This leads to a contradiction. Thus, there exists $h$ such that the number of vertices in $T_{h}$ with degree no less than $\frac{1}{10}\left|X_{i}Y_{j}\right|$ is at least $\frac{n}{8N}$. We claim that there are at least $\frac{1}{100}|X_{i}||Y_{j}|$ elements in $X_{i}Y_{j}$ such that their codegree in $T_{h}$ is no less than $\sqrt{n}$. Otherwise, a contradiction holds:
$\frac{n}{8N}\cdot\frac{1}{10}\left|X_{i}Y_{j}\right|$ $\leq E(X_{i}Y_{j},T_{h})$
$\leq\sqrt{n}\cdot\left|X_{i}Y_{j}\right|+\frac{n}{N}\cdot\frac{1}{100}\left|X_{i}Y_{j}\right|$
$<\frac{n}{90N}\left|X_{i}Y_{j}\right|.$
Now consider a bipartite graph $B^{\prime}$ with vertex set $V(X_{i},Y_{j})$, and its edge density is no less than $\frac{1}{100}$. We claim that there are $0.001|X_{i}|$ vertices in $X_{i}$ such that their degree in $B^{\prime}$ is no less than $0.001|Y_{j}|$. Otherwise, the edge density of $B^{\prime}$ is no more than $1\times 0.001+0.001\times 1<\frac{1}{100}$, which leads to a contradiction. Let these vertices be $X^{0}_{i}$ then the claim holds. ∎
To ensure consistency in the number of points in subsequent work, we should rearrange the order of every triple to be even.
Claim 4.9.
We can embed at most $N/2$ cycles of length greater than $3$ using only the vertices from the triples such that every triple minus these cycles has an even number of vertices.
Proof.
Since $N$ and $\sum_{i\in N}|T_{i}|$ are even, the number of triples with an odd order is also even. We use the following progress until there is no odd order triple. If there exists $T_{i}$ and $T_{j}$ having odd orders, using Claim 4.8, there exist a vertex $x$ in $T_{i}$, a subset of vertices $Y^{\prime}_{j}\subseteq Y_{j}$ with order no less than $0.001|Y_{j}|$ as well as a triple $T_{p}$, such that for any $y\in Y^{\prime}_{j}$, there exist $\sqrt{n}$ vertices $t_{j}\in T_{p}$ such that $(xt_{j}y_{j})\in E(\mathcal{H})$.
Applying Lemma 3.8, we can construct a loose path with both of its endpoints $y_{1},y_{2}$ in $Y_{j}$ of a certain length (more than $2$) and there exists $t_{1},t_{2}\in T_{p}$ (since $\sqrt{n}\gg(N/2)\log n$, there are always sufficiently many such vertices available) such that $(xt_{1}y_{1}),(xt_{2}y_{2})\in E(\mathcal{H})$. This path, along with the vertices $x$ and 2 vertices $t_{1},t_{2}$ in $T_{p}$, form a loose cycle $C_{ij}$ of length more than $3$ in $\mathcal{C}$. We then remove this cycle from $\mathcal{C}$ and delete the vertices that were used in this process.
During this process, $|T_{i}|$ decreased by $1$, $|T_{j}|$ decreased by $|C_{ij}-3|$ and $|T_{p}|$ decreased by $2$, and the other triple’s order remained unchanged, thus the parity of all triples remains unchanged while $|T_{i}|$ and $|T_{j}|$ changed from an odd number to an even number. By this progress no more than $N/2$ times all $|T_{i}|$ turns even. ∎
Without loss of generality, remove all vertices that have been used for embedding cycles from the triples. Still, let $T_{i}=(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ denote the triple for every $i\in[N]$. The remaining $r^{\prime}$ vertex-disjoint loose cycles are denoted by $C_{n_{1}},C_{n_{2}},\dots,C_{n_{r^{\prime}}}$, where $k^{\prime}$ of these cycles are odd. For each $i$, we have $n_{i}<\log n$, and the sequence $n_{1},n_{2},\dots,n_{r^{\prime}}$ is listed in increasing order, i.e., $n_{1}\leq n_{2}\leq\dots\leq n_{r^{\prime}}$.
Define $\mathcal{C}_{0}=\cup_{i=1}^{r^{\prime}}C_{i}$, $\mathcal{C}_{1}=\cup_{i=1}^{r^{\prime}-2(N-1)}C_{i}$, and $\mathcal{C}_{2}=\cup_{i=r^{\prime}+1-2(N-1)}^{r^{\prime}}C_{i}$ as the disjoint unions of loose cycles. Let $C_{lgc}$ denote the longest cycle in $\mathcal{C}_{1}$. Denote the number of cycles in $\mathcal{C}_{1}$ by $r^{\prime}_{1}$, and let $k^{\prime}_{1}$ be the number of odd cycles in $\mathcal{C}_{1}$.
Claim 4.10 (Pre-partition).
There exists a partition $\mathcal{C}_{1}=\cup_{i=1}^{N}\overline{\mathcal{F}}_{i}$ with ¯$\overline{\mathcal{F}}_{i}\cap\overline{\mathcal{F}}_{j}=\emptyset$ when $i\neq j$ such that for any $i\in[N]$, $\big{|}\sum_{j\in\overline{\mathcal{F}}_{i}}|C_{j}|-|T_{i}|\big{|}\leq N^{2}\log n$. Moreover, the number of odd cycles in any ¯$\overline{\mathcal{F}}_{i}$ is no more than $\frac{k(1+\eta^{3.2})}{N}$.
Proof.
Without loss of generality, we prove a stronger version of the result. Note that for any two triples, the difference in their orders is no more than $\eta^{4}m+N\log n+4n\leq\eta^{3.5}m$, and for each pair of cycles, the difference in their orders is bounded by $|C_{lgc}|\leq\log n$. Given a partition $\mathcal{P}:\mathcal{C}_{1}=\cup_{i=1}^{N}\overline{\mathcal{F}}_{i}$, define $\phi_{i}$ as the number of odd cycles in ¯$\overline{\mathcal{F}}_{i}$, and define $\beta_{i}$ as the number of even cycles in ¯$\overline{\mathcal{F}}_{i}$. By Lemma A.4 (with $k^{\prime}_{1}=q$, $r^{\prime}_{1}-k^{\prime}_{1}=q$, and $a_{i}=|T_{i}|/\sum|T_{i}|$), there exists at least one partition such that $\phi_{i}\in\left[k^{\prime}_{1}|T_{i}|/\sum|T_{i}|-1,k^{\prime}_{1}|T_{i}|/\sum|T_{i}|+1\right]$ and $\beta_{i}\in\left[(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|-1,(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|+1\right]$ for every $i\in[N]$.
We claim that such a partition exists, satisfying $\phi_{i}\in\left[k^{\prime}_{1}|T_{i}|/\sum|T_{i}|-1,k^{\prime}_{1}|T_{i}|/\sum|T_{i}|+1\right]$ and $\beta_{i}\in\left[(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|-1,(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|+1\right]$ for every $i\in[N]$. Furthermore, for any $i\in[N]$, we have $\left|\sum_{j\in\overline{\mathcal{F}}_{i}}|C_{j}|-|T_{i}|\right|\leq N^{2}\log n$. Otherwise, a contradiction arises under the following two cases. This implies that the number of odd cycles in any ¯$\overline{\mathcal{F}}_{i}$ is at most $k^{\prime}_{1}|T_{i}|/\sum|T_{i}|+1\leq\frac{k(1+\eta^{3.2})}{N}$.
Case 1: For any such partition $\mathcal{P}$ we say $i$ is 1-bad if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|>N^{2}\log n$, and let $\mathcal{S}=\sum_{i\ \text{is}\ 1\text{-bad}}\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\big{)}$. If for any such partition with $\phi_{i}\in[k^{\prime}_{1}|T_{i}|/\sum|T_{i}|-1,k^{\prime}_{1}|T_{i}|/\sum|T_{i}|+1]$ and $\beta_{i}\in[(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|-1,(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|+1]$ there exists $i\in[N]$ such that $i$ is 1-bad, choose one such partition with minimum $\mathcal{S}$. Note that there must be $i^{\prime}\in[N]$ with $\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|-|T_{i^{\prime}}|<-2\log n$ otherwise $\sum_{C\in\mathcal{C}_{1}}|C|-\sum|T_{i}|>(N^{2}-2N+2)\log n\geq\sum_{C\in\mathcal{C}_{0}}|C|-\sum|T_{i}|$, which leads to a contradiction. Thus $\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|+2|C_{lgc}|\big{)}/|T_{i^{\prime}}|\leq 1.$
We claim that there exists $C_{x_{i}}\in\overline{\mathcal{F}}_{i}$ and $C_{x_{i^{\prime}}}\in\overline{\mathcal{F}}_{i^{\prime}}$ such that $|C_{x_{i}}|>|C_{x_{i^{\prime}}}|$ and $C_{x_{i}},C_{x_{i^{\prime}}}$ have same parity. Let $C^{i}_{lgoc}$ be the longest odd cycle in ¯$\overline{\mathcal{F}}_{i}$ and $C^{i}_{lgec}$ be the longest even cycle in ¯$\overline{\mathcal{F}}_{i}$. Otherwise since $(\phi_{i}-1)/(\phi_{i^{\prime}}+1)\leq|T_{i}|/|T_{i^{\prime}}|$ and $(\beta_{i}-1)/(\beta_{i^{\prime}}+1)\leq|T_{i}|/|T_{i^{\prime}}|$, thus
$\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-2|C_{lgc}|\big{)}/\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|+2|C_{lgc}|\big{)}$
$\leq\big{(}(\phi_{i}-1)\cdot|C^{i}_{lgoc}|+(\beta_{i}-1)\cdot|C^{i}_{lgec}|\big{)}/\big{(}(\phi_{i^{\prime}}+1)\cdot|C^{i}_{lgoc}|+(\beta_{i^{\prime}}+1)\cdot|C^{i}_{lgec}|\big{)}$
$\leq|T_{i}|/|T_{i^{\prime}}|.$
Hence, $\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-2|C_{lgc}|\big{)}/|T_{i}|\leq\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|+2|C_{lgc}|\big{)}/|T_{i^{\prime}}|\leq 1$, which implies $N^{2}\log n<\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\leq 2|C_{lgc}|$, and this leads to a contradiction. Thus, delete $C_{x_{i}}$ from ¯$\overline{\mathcal{F}}_{i}$ and add it into ¯$\overline{\mathcal{F}}_{i^{\prime}}$ while delete $C_{x_{i^{\prime}}}$ from ¯$\overline{\mathcal{F}}_{i^{\prime}}$ and add it into ¯$\overline{\mathcal{F}}_{i}$ makes $\mathcal{S}$ decrease by at least $2$, and this leads to a contradiction with the minimality of $\mathcal{S}$.
Case 2: For any such partition $\mathcal{P}$ we say $i$ is 2-bad if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|<-N^{2}\log n$, and let $\mathcal{S^{\prime}}=\sum_{i\ \text{is}\ 2\text{-bad}}\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\big{)}$. If for any such partition with $\phi_{i}\in[k^{\prime}_{1}|T_{i}|/\sum|T_{i}|-1,k^{\prime}_{1}|T_{i}|/\sum|T_{i}|+1]$ and $\beta_{i}\in[(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|-1,(r^{\prime}_{1}-k^{\prime}_{1})|T_{i}|/\sum|T_{i}|+1]$ there exists $i\in[N]$ such that $i$ is 2-bad, choose one such partition with maximum $\mathcal{S^{\prime}}$. Note that there must be $i^{\prime}\in[N]$ with $\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|-|T_{i^{\prime}}|>2\log n$ otherwise $\sum_{C\in\mathcal{C}_{1}}|C|-\sum|T_{i}|<-(N^{2}-2N+2)\log n\leq-2N\log n<-\sum_{C\in\mathcal{C}_{2}}|C|$, thus $\sum_{C\in\mathcal{C}_{0}}|C|-\sum|T_{i}|<0$, which leads to a contradiction. Thus $\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|-2|C_{lgc}|\big{)}/|T_{i^{\prime}}|\geq 1.$
We claim that there exists $C_{x_{i}}\in\overline{\mathcal{F}}_{i}$ and $C_{x_{i^{\prime}}}\in\overline{\mathcal{F}}_{i^{\prime}}$ such that $|C_{x_{i}}|<|C_{x_{i^{\prime}}}|$ and $C_{x_{i}},C_{x_{i^{\prime}}}$ have the same parity. Let $C^{i}_{stoc}$ be the shortest odd cycle in ¯$\overline{\mathcal{F}}_{i}$ and $C^{i}_{stec}$ be the shortest even cycle in ¯$\overline{\mathcal{F}}_{i}$. Otherwise, since $(\phi_{i}+1)/(\phi_{i^{\prime}}-1)\geq|T_{i}|/|T_{i^{\prime}}|$ and $(\beta_{i}+1)/(\beta_{i^{\prime}}-1)\geq|T_{i}|/|T_{i^{\prime}}|$, thus
$\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|+2|C_{lgc}|\big{)}/\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|-2|C_{lgc}|\big{)}$
$\geq\big{(}(\phi_{i}+1)\cdot|C^{i}_{stoc}|+(\beta_{i}+1)\cdot|C^{i}_{stec}|\big{)}/\big{(}(\phi_{i^{\prime}}-1)\cdot|C^{i}_{stoc}|+(\beta_{i^{\prime}}-1)\cdot|C^{i}_{stec}|\big{)}$
$\geq|T_{i}|/|T_{i^{\prime}}|.$
Hence, $\big{(}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|+2|C_{lgc}|\big{)}/|T_{i}|\geq\big{(}\sum_{C\in\overline{\mathcal{F}}_{i^{\prime}}}|C|-2|C_{lgc}|\big{)}/|T_{i^{\prime}}|\geq 1$, which implies $-2|C_{lgc}|\leq\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|<-N^{2}\log n$, this leads to a contradiction. Thus delete $C_{x_{i}}$ from ¯$\overline{\mathcal{F}}_{i}$ and add it into ¯$\overline{\mathcal{F}}_{i^{\prime}}$ while delete $C_{x_{i^{\prime}}}$ from ¯$\overline{\mathcal{F}}_{i^{\prime}}$ and add it into ¯$\overline{\mathcal{F}}_{i}$ makes $\mathcal{S^{\prime}}$ increase by at least $2$, and moreover since $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|<-N^{2}\log n+\log n<0$, by this step $\mathcal{S}$ doesn’t increase. This leads to a contradiction with the maximality of $\mathcal{S^{\prime}}$. ∎
Claim 4.11.
If $|C_{lgc}|\geq 16$, then there exists an embedding $\Psi$ from $\mathcal{C}_{2}$ into $\cup^{N}_{i=1}T_{i}$ such that the following holds:
We can find a partition $\mathcal{C}_{1}=\cup_{i=1}^{N}\mathcal{F}_{i}$ with $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset$ when $i\neq j$ such that for any $i\in[N]$, $|V(T_{i}\setminus\cup_{C\in\mathcal{C}_{2}}\Psi(V(C))|=\sum_{C\in\mathcal{F}_{i}}|C|$. Moreover, the number of odd cycles in any $\mathcal{F}_{i}$ is no more than $\frac{k(1+\eta^{3})}{N}$.
If $|C_{lgc}|<16$, there exists a partition $\mathcal{C}_{0}=\cup_{i=1}^{N}\mathcal{F}_{i}$ with $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset$ when $i\neq j$ such that for any $i\in[N]$, $|V(T_{i})|=\sum_{C\in\mathcal{F}_{i}}|C|$. Moreover, the number of odd cycles in any $\mathcal{F}_{i}$ is no more than $\frac{k(1+\eta^{3})}{N}$.
Proof.
If $|C_{lgc}|\geq 16:$
By Claim 4.10, there exists a partition $\mathcal{C}_{1}=\cup_{i=1}^{N}\overline{\mathcal{F}}_{i}$ with ¯$\overline{\mathcal{F}}_{i}\cap\overline{\mathcal{F}}_{j}=\emptyset$ when $i\neq j$ such that for any $i\in[N]$, $\big{|}\sum_{j\in\overline{\mathcal{F}}_{i}}|C_{j}|-|T_{i}|\big{|}\leq N^{2}\log n$. Moreover, the number of odd cycles in any ¯$\overline{\mathcal{F}}_{i}$ is no more than $\frac{k(1+\eta^{3.2})}{N}$.
For any $i\in[N-1]$, by Claim 4.8, there exists a triple $T_{h}^{l}$ (and $T_{h}^{l^{\prime}}$) with $l=l(i),l^{\prime}=l^{\prime}(i)$, and subsets $X_{i}^{0}\subseteq X_{i}$ (and $Y_{i}^{0}\subseteq Y_{i}$) where $|X_{i}^{0}|\geq 0.001|X_{i}|$ (and $|Y_{i}^{0}|\geq 0.001|Y_{i}|$), such that for any $x_{i}\in X_{i}^{0}$ (and $y_{i}\in Y_{i}^{0}$), there exist at least $0.001|Y_{i+1}|$ (and $0.001|X_{i+1}|$) vertices $y_{i+1}\in Y_{i+1}$ (and $x_{i+1}\in X_{i+1}$) such that for each $y_{i+1}$ (and $x_{i+1}$), there are at least $\sqrt{n}$ values of $t$ in $T_{h}^{l}$ (and $T_{h}^{l^{\prime}}$) with $(x_{i}ty_{i+1})\in E(\mathcal{H})$ (and $(x_{i+1}ty_{i})\in E(\mathcal{H})$). Since $\sqrt{n}\gg 2N\log n$, there are always sufficiently many such vertices $t$ available. For each triple, record the number of times it appears in this form. In the subsequent proof, assume that the count for this triple is reduced by subtracting $2$ times this recorded number.
We use the following cyclic process for each $i\in[N-1]$:
In step $i$, remove some cycles from ¯$\overline{\mathcal{F}}_{i}$ to ¯$\overline{\mathcal{F}}_{i+1}$ if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\geq-6$ or remove some cycles from ¯$\overline{\mathcal{F}}_{i+1}$ to ¯$\overline{\mathcal{F}}_{i}$ if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\leq-|C_{lgc}|-6$ such that the sum of their orders is no less than $|T_{i}|-|C_{lgc}|-6$ and no more than $|T_{i}|-6$. Denote these cycles as $\mathcal{F}_{i}$.
Since $\big{|}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\big{|}\leq\max\{N^{2}\log n,i\cdot N^{3}\log n\}=i\cdot N^{3}\log n$, by adding/deleting no more than $\frac{i}{|C_{6}|}\cdot N^{3}\log n+|C_{lgc}|+6\leq\frac{i}{5}\cdot N^{3}\log n$ cycles into/from ¯$\overline{\mathcal{F}}_{i}$ this progress could be reached. Thus the number of odd cycles in $\mathcal{F}_{i}$ is no more than
$\frac{k(1+\eta^{3.2})}{N}+(i-1)\cdot N^{3}\log n+\frac{i}{5}\cdot N^{3}\log n\leq\frac{k(1+\eta^{3})}{N}.$
We embed 2 paths $L^{\prime}$ and $L^{\prime\prime}$ ($|L^{\prime}|,|L^{\prime\prime}|\geq 3$) in $T_{i}$ with its end vertices $x^{\prime},x^{\prime\prime}$ in $X_{i}^{0}\cap T_{i}$ and $y^{\prime},y^{\prime\prime}$ in $Y_{i}^{0}\cap T_{i}$ such that $|L^{\prime}|,|L^{\prime\prime}|=\big{|}|T_{i}|-\sum_{C\in\mathcal{F}_{i}}|C|\big{|}/2$ by Lemma 3.8. For both of $L^{\prime}$’s end vertices $x^{\prime}$ and $y^{\prime}$, there exists $Y^{0}_{i+1}\subseteq Y_{i+1}$ with $|Y^{0}_{i+1}|\geq 0.001|Y_{i+1}|$ and $X^{0}_{i+1}\subseteq X_{i+1}$ with $|X^{0}_{i+1}|\geq 0.001|X_{i+1}|$ such that for every vertex $y^{0}_{i+1}$ in $Y^{0}_{i+1}$(and $x^{0}_{i+1}$ in $X^{0}_{i+1}$), there exist $l=l(i),l^{\prime}=l^{\prime}(i)$ and at least $\sqrt{n}$ numbers of $t$ in $T_{h}^{l}$ (and $T_{h}^{l^{\prime}}$) with $(x^{\prime}ty^{0}_{i+1})\in E(\mathcal{H})$ (and $(x^{0}_{i+1}ty^{\prime})\in E(\mathcal{H})$). By Lemma 3.8, embed a path $L_{0}^{\prime}$ (where $|L_{0}^{\prime}|\geq 3$) in $T_{i+1}$, with its end vertices $y_{0}^{\prime}$ and $x_{0}^{\prime}$ in $Y^{0}_{i+1}$ and $X^{0}_{i+1}$, respectively, such that $|L^{\prime}|+|L_{0}^{\prime}|+2$ equals $|C^{\prime}|$ for some $C^{\prime}\in\mathcal{C}_{2}$. By Claim 4.8, there exists $t^{\prime}_{1}\in T_{h}^{l}$ and $t^{\prime}_{2}\in T_{h}^{l^{\prime}}$ such that $(x^{\prime}y_{0}^{\prime}t^{\prime}_{1}),(y^{\prime}x_{0}^{\prime}t^{\prime}_{2})\in E(\mathcal{H})$. Thus, $L^{\prime},L^{\prime}_{0}$ together with 2 vertices $t^{\prime}_{1},t^{\prime}_{2}$ form a cycle $C^{\prime}$ in $\mathcal{C}_{2}$, then delete $C^{\prime}$ in $\mathcal{C}_{2}$. Denote $\Psi(V(C^{\prime}))$ to be the vertices in $L^{\prime},L^{\prime}_{0}$ and $t^{\prime}_{1},t^{\prime}_{2}$, then remove every vertices in $\Psi(V(C^{\prime}))$ from $V(\cup^{N}_{i=1}T_{i})$. Then do a same progress for $L^{\prime\prime}$ such that $L^{\prime\prime}$ together with a path $L^{\prime\prime}_{0}$ ($|L_{0}^{\prime\prime}|\geq 3$) embedded in $T_{i+1}$ and 2 vertices in $T_{h}^{l}$ and $T_{h}^{l^{\prime}}$ form a cycle $C^{\prime\prime}$ in $\mathcal{C}_{2}$, then delete $C^{\prime\prime}$ in $\mathcal{C}_{2}$. Denote $\Psi(V(C^{\prime\prime}))$ to be the vertices in $L^{\prime\prime},L^{\prime\prime}_{0}$ and $t^{\prime\prime}_{1},t^{\prime\prime}_{2}$, then remove every vertices in $\Psi(V(C^{\prime\prime}))$ from $V(\cup^{N}_{i=1}T_{i})$.
These cycles exist since for any cycle $C_{x}$ in $\mathcal{C}_{2}$, $|C_{x}|\geq 16$, thus $62|C_{x}|-10\geq|C_{lgc}|+$ 6, and this implies that for any even number $w$ in $[|T_{i}|-|C_{lgc}|-6,|T_{i}|-6]$, $w/2\geq 3$ and $|C_{x}|-w/2\geq|C_{x}|-(|C_{lgc}|+6)/2\geq 5$, hence $|L^{\prime}|,|L^{\prime\prime}|,|L_{0}^{\prime}|,|L^{\prime\prime}_{0}|\geq 3$ holds.
Having done this step, we have
$\big{|}\sum_{C\in\overline{\mathcal{F}}_{i+1}}|C|-|T_{i+1}|\big{|}$
$\leq i\cdot N^{3}\log n+2\log n+|C_{lgc}|+6+N^{2}\log n$
$\leq(i+1)\cdot N^{3}\log n,$
and the number of odd cycles in ¯$\overline{\mathcal{F}}_{i+1}$ is no more than $\frac{k(1+\eta^{3.2})}{N}+\frac{i}{5}\cdot N^{3}\log n\leq\frac{k(1+\eta^{3.2})}{N}+i\cdot N^{3}\log n.$ This implies the following Fact:
Fact 4.12.
Having finished step $i$ we have $\big{|}\sum_{C\in\overline{\mathcal{F}}_{i+1}}|C|-|T_{i+1}|\big{|}\leq(i+1)\cdot N^{3}\log n$ and the number of odd cycles in ¯$\overline{\mathcal{F}}_{i+1}$ is no more than $\frac{k(1+\eta^{3.2})}{N}+i\cdot N^{3}\log n$.
When it comes to $\mathcal{F}_{N}$, since every other triple is used up, its order is the union of the cycles left in $\mathcal{C}_{1}$. And the number of odd cycles in it is no more than $\frac{k(1+\eta^{3.2})}{N}+N^{4}\log n\leq\frac{k(1+\eta^{3})}{N}$.
If $|C_{lgc}|<16$:
Let $c_{x}$ denote the number of cycles with order $x$ and $c^{i}_{x}$ denote the number of cycles with order $x$ in ¯$\overline{\mathcal{F}}_{i}$. Partition $\mathcal{C}_{0}=\cup_{i=1}^{N}\overline{\mathcal{F}}_{i}$ with ¯$\overline{\mathcal{F}}_{i}\cap\overline{\mathcal{F}}_{j}=\emptyset$ when $i\neq j$ such that for any $i\in[N]$ and any $x$, $c^{i}_{x}$ is no more than $c_{x}|T_{i}|/\sum|T_{i}|+1$ and no less than $c_{x}|T_{i}|/\sum|T_{i}|-1$ for any $i\in[N]$ and every $x$. The existence of this partition is ensured by Lemma A.4. Note that for any $i\in[N]$ the number of odd cycles in ¯$\overline{\mathcal{F}}_{i}$ is no more than $k^{\prime}|T_{i}|/\sum|T_{i}|+\sum_{x}1\leq k^{\prime}|T_{i}|/\sum|T_{i}|+3N.$ Let $A\{a_{1},a_{2},\dots\}\subseteq\{6,8,10,12,14\}$ denote the $x$ such that $c_{x}\geq n\eta^{3}/2x$. If $A$ is $\{8,12\}$ or any set with order $1$, then the problem could be solved by Lemma 4.2 and Lemma 4.3, Thus, we only need to consider the case where $A$ is neither $\{8,12\}$ nor any set with order $1$. In this case, the greatest common divisor of $a_{1},a_{2},\dots$ is $2$, therefore there exist integers $b_{1},b_{2},\dots$ such that $a_{1}b_{1}+a_{2}b_{2}+\dots=2$.
Define a A-transformation from a set of cycles $\mathcal{C^{\prime\prime}}$ to another set of cycles $\mathcal{C^{\prime}}$: $\mathcal{A}:\mathcal{C^{\prime\prime}}\rightarrow\mathcal{C^{\prime}}$ as following: for every $a_{l}\in A$, if $b_{l}$ is positive, then remove $b_{l}$ numbers of $C_{a_{l}}$ from $\mathcal{C^{\prime\prime}}$ into $\mathcal{C^{\prime}}$; if $b_{l}$ is negative, then remove $-b_{l}$ numbers of $C_{a_{l}}$ from $\mathcal{C^{\prime}}$ into $\mathcal{C^{\prime\prime}}$. After a A-transformation from $\mathcal{C^{\prime\prime}}$ to $\mathcal{C^{\prime}}$, $\sum_{C\in\mathcal{C^{\prime}}}|C|$ increases by 2 compared to before while $\sum_{C\in\mathcal{C^{\prime\prime}}}|C|$ decreases by 2 compared to before.
Note that for every $i$ and every $x$, $\big{|}x\cdot c_{x}^{i}-x\cdot c_{x}|T_{i}|/\sum|T_{i}|\big{|}\leq x$, thus we have:
$\big{|}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\big{|}$
$\leq\sum_{x}\big{|}x\cdot c_{x}^{i}-x\cdot c_{x}|T_{i}|/\sum|T_{i}|\big{|}$
$\leq 6+8+10+12+14+2(N-1)\log n\leq 2N\log n.$
We use the following cyclic process for each $i\in[N-1]$:
In step $i$, do some A-transformation from ¯$\overline{\mathcal{F}}_{i}$ to ¯$\overline{\mathcal{F}}_{i+1}$ if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|>0$ or do some A-transformation from ¯$\overline{\mathcal{F}}_{i+1}$ to ¯$\overline{\mathcal{F}}_{i}$ if $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|<0$ until $\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|=0$. Since $\big{|}\sum_{C\in\overline{\mathcal{F}}_{i}}|C|-|T_{i}|\big{|}\leq i\cdot 2N\log n$, by no more than $i\cdot N\log n$ numbers of A-transformation this progress could be reached. Since every $x\in A$ satisfies $c_{x}\geq n\eta^{3}/2x$ and thus for every $x\in A$ and every $j\in[N]$, $c^{j}_{x}\gg i\cdot N\log n\cdot(|b_{1}|+|b_{2}|+\dots)$, these A-transformation exists.
Denote these cycles as $\mathcal{F}_{i}$. Note that the number of odd cycles in $\mathcal{F}_{i}$ is no more than $k^{\prime}|T_{i}|/\sum|T_{i}|+3N+(i-1)\cdot N\log n(|b_{1}|+|b_{2}|+\dots)+i\cdot N\log n(|b_{1}|+|b_{2}|+\dots)\leq\frac{k(1+\eta^{3})}{N}$.
Having done this step, we have
$\big{|}\sum_{C\in\overline{\mathcal{F}}_{i+1}}|C|-|T_{i+1}|\big{|}\leq 2N\log n+i\cdot 2N\log n=(i+1)\cdot 2N\log n,$
and the number of odd cycles in ¯$\overline{\mathcal{F}}_{i+1}$ is no more than $k^{\prime}|T_{i}|/\sum|T_{i}|+3N+i\cdot N\log n(|b_{1}|+|b_{2}|+\dots).$ This implies the following Fact:
Fact 4.13.
Having finished step $i$ we have $\big{|}\sum_{C\in\overline{\mathcal{F}}_{i+1}}|C|-|T_{i+1}|\big{|}\leq(i+1)\cdot 2N\log n$ and the number of odd cycles in ¯$\overline{\mathcal{F}}_{i+1}$ is no more than $k^{\prime}|T_{i+1}|/\sum|T_{i+1}|+3N+i\cdot N\log n(|b_{1}|+|b_{2}|+\dots)$.
When it comes to $\mathcal{F}_{N}$, since every other triple is used up, its order is the union of the cycles left in $\mathcal{C}_{0}$. And the number of odd cycles in it is no more than $k^{\prime}|T_{i}|/\sum|T_{i}|+3N+N^{2}\log n(|b_{1}|+|b_{2}|+\dots)\leq\frac{k(1+\eta^{3})}{N}$.
4.3 Step 3
Now, for each triple still denoted as $T_{i}=(X_{i},Y_{i},Z_{i})_{\mathcal{H^{\prime}}}$ for every $i$, since no more than $N^{10}\log n$ vertices were used during the order adjustment to make the triples even and embed the cycles in $\mathcal{C}_{2}$, it remains weakly $(2\varepsilon^{*},d^{*}/2)$-half-superregular by Lemma 3.1. Moreover for every $i\in[N]$ we have:
$(2a+b)/a\cdot\min\{|X_{i}|,|Y_{i}|\}-(|X_{i}|+|Y_{i}|)-|Z_{i}|\geq\sqrt{m}-(a+b)N^{10}\log n>0.$
And $(|Z_{i}|+\sqrt{m})/(|X_{i}|+|Y_{i}|+2N^{10}\log n)\leq b/2a$, thus
$|Z_{i}|/(|X_{i}|+|Y_{i}|)<(|Z_{i}|+\sqrt{m})/(|X_{i}|+|Y_{i}|+2N^{10}\log n)\leq b/2a.$
Also we have $|Z_{i}|-\max\{|X_{i}|,|Y_{i}|\}\geq\eta^{3}n-N^{10}\log n>0.$ Therefore, if we represent $(X_{i},Y_{i},Z_{i})$ as $(V_{1},V_{2},V_{3})$ with $|V_{1}|\leq|V_{2}|\leq|V_{3}|$, we can conclude that:
$\frac{|V_{1}|+|V_{2}|}{|V_{3}|}\geq\frac{2a}{b};$
$\frac{a+b}{a}|V_{1}|\geq|V_{2}|+|V_{3}|.$
Claim 4.14.
For every $i\in[N]$, there is a perfect embedding from $\mathcal{F}_{i}$ into $T_{i}$ by Lemma 3.7.
Proof.
We prove it for each $i\in[N]$. Denote $|T_{i}|=m^{*}$ and let the number of odd cycles in $\mathcal{F}_{i}$ be $k^{\prime}_{i}$. It follows that $0<\frac{1}{m^{*}}\ll\varepsilon^{*}\ll d^{*},\xi\ll\eta\ll 1$ and $T_{i}$ is weakly $(2\varepsilon^{*},d^{*}/2)$-half-superregular with $m^{*}=\sum_{C\in\mathcal{F}_{i}}|C|$.
Since $k^{\prime}_{i}\leq\frac{k(1+\eta^{3})}{N}$ and $\frac{(1-3\sqrt{\xi})n}{N}\leq m^{*}\leq\frac{n}{N}$, we have
$\frac{2m^{*}-4k^{\prime}_{i}}{m^{*}+2k^{\prime}_{i}}\geq\frac{2n(1-3\sqrt{\xi})-4k(1+\eta^{3})}{(n+2k)(1+\eta^{3})}\geq\frac{2n-4k}{(n+2k)(1+\eta^{2.5})}.$
Since $a,b$ are $(n,k,\eta)$-good, we have $b/a\leq 2-\eta^{2}$ and
$\frac{b}{a}\leq\frac{2n-4k}{n+2k}-\eta^{2}\leq(1+\eta^{2.5})\frac{2m^{*}-4k^{\prime}_{i}}{m^{*}+2k^{\prime}_{i}}-\eta^{2}\leq\frac{2m^{*}-4k^{\prime}_{i}}{m^{*}+2k^{\prime}_{i}}-0.1\eta^{2}.$
Thus, we obtain:
$\frac{m^{*}a}{2a+b}=\frac{m^{*}}{2+b/a}\geq m^{*}/4+k^{\prime}/2.$
Let $2x_{i}=|V_{j}|-x_{j}+|V_{h}|-x_{h}$ for every $\{i,j,h\}=\{1,2,3\}$. We have:
$x_{1}=\frac{1}{2}(|V_{2}|+|V_{3}|-|V_{1}|),\quad x_{2}=\frac{1}{2}(|V_{1}|+|V_{3}|-|V_{2}|),\quad x_{3}=\frac{1}{2}(|V_{1}|+|V_{2}|-|V_{3}|).$
We claim that there exists a tripartite graph $G_{0}$ with each part of order $x_{1},x_{2},x_{3}$, consisting of vertex-disjoint cycles $C_{\frac{m_{1}}{2}},C_{\frac{m_{2}}{2}},\ldots,C_{\frac{m_{r}}{2}}$ where $k^{\prime}_{i}$ of these are odd. Let $\{i,j,h\}=\{1,2,3\}$ and $x_{h}=\max\{x_{i},x_{j},x_{h}\}$. Since $\frac{a+b}{a}|V_{1}|\geq|V_{2}|+|V_{3}|=m^{*}-|V_{1}|$, we have $|V_{1}|\geq\frac{m^{*}a}{2a+b}$. Thus the minimum degree of the complete graph $K_{x_{i}x_{j}x_{h}}$ is no less than:
$\min\{x_{i}+x_{j},x_{j}+x_{h},x_{i}+x_{h}\}$
$=|V_{1}|\geq\frac{m^{*}a}{2a+b}\geq m^{*}/4+k^{\prime}_{i}/2$
$=\frac{x_{1}+x_{2}+x_{3}}{2}+k^{\prime}_{i}/2.$
By Lemma 3.11, there exists a $3$-partite graph $G_{0}$ with each part of order $x_{1},x_{2},x_{3}$, consisting of vertex-disjoint cycles $C_{\frac{m_{1}}{2}},C_{\frac{m_{2}}{2}},\ldots,C_{\frac{m_{r}}{2}}$ where $k^{\prime}_{i}$ of these are odd.
Taking a uniformly random partition of $V_{1},V_{2},V_{3}$ into disjoint sets such that $V_{1}=X_{1}\cup X_{2}$, $V_{2}=Y_{1}\cup Y_{2}$, $V_{3}=Z_{1}\cup Z_{2}$ with $|X_{1}|=x_{1}$, $|Y_{1}|=x_{2}$ and $|Z_{1}|=x_{3}$. For any $i\in[3]$, the number of edges in $G_{0}$ adjacent to the $x_{i}$ part is $2x_{i}$, and for any $i,j,h\in[3]$, the number of edges between the $x_{i}$ part and the $x_{j}$ part is $x_{i}+x_{j}-x_{h}$, and $(x_{i}+x_{j}-x_{h})+x_{h}=|V_{h}|$. Since $b/a\leq 2-\eta^{2}$ and $\frac{a+b}{a}|V_{1}|\geq|V_{2}|+|V_{3}|$,
$3|V_{1}|-|V_{2}|-|V_{3}|\geq 3|V_{1}|-\frac{a+b}{a}|V_{1}|\geq m^{*}\frac{a}{2a+b}(3-\frac{a+b}{a})$
$=m^{*}\frac{2a-b}{2a+b}\geq\frac{\eta^{2}m^{*}}{4},$
we have $|V_{i}|-x_{i}=1.5|V_{i}|-0.5(|V_{j}|+|V_{h}|)\geq\frac{\eta^{2}m^{*}}{8}$, thus $\min\{|X_{2}|,|Y_{2}|,|Z_{2}|\}\geq\frac{\eta^{2}m^{*}}{8}$. On another aspect, since $\frac{|V_{1}|+|V_{2}|}{|V_{3}|}\geq 2a/b$, we have:
$\min\{|X_{1}|,|Y_{1}|,|Z_{1}|\}\geq\frac{1}{2}(|V_{1}|+|V_{2}|-|V_{3}|)$
$\geq\frac{1}{2}m^{*}\frac{2a-b}{2a+b}\geq\frac{\eta^{2}m^{*}}{8},$
thus $\min\{|X_{1}|,|Y_{1}|,|Z_{1}|\}\geq\frac{\eta^{2}m^{*}}{8}$.
By Lemma 3.1, we can assert that, with high probability, each of the triples $(X_{2},Y_{1},Z_{2})_{\mathcal{H^{\prime}}}$, $(X_{2},Z_{1},Y_{2})_{\mathcal{H^{\prime}}}$, and $(Z_{2},X_{1},Y_{2})_{\mathcal{H^{\prime}}}$ is $(\frac{8}{\eta^{2}}\varepsilon^{*},{d^{*}}^{2}/64)$-half-superregular, thus by union bound there exists such a partition such that all of the triples $(X_{2},Y_{1},Z_{2})_{\mathcal{H^{\prime}}}$, $(X_{2},Z_{1},Y_{2})_{\mathcal{H^{\prime}}}$, and $(Z_{2},X_{1},Y_{2})_{\mathcal{H^{\prime}}}$ are $(\frac{8}{\eta^{2}}\varepsilon^{*},{d^{*}}^{2}/64)$-half-superregular, and let $G_{0}=G_{0}(X_{1},Y_{1},Z_{1})$. Applying Theorem 3.7 with $r=3$, we can embed a 1-expansion of $G_{0}$, which contains all of the cycles in $\mathcal{F}_{i}$.
5 Concluding Remarks
  • In this paper, we employed the regularity method and the blow-up lemma to establish an asymptotically tight bound for the embedding problem of disjoint loose cycles in 3-graphs. Based on our findings, we propose the following conjecture regarding the tight bounds:
    Conjecture 5.1.
    There exists an integer $n_{0}$ such that for any $n\geq n_{0}$, the following holds. Let $\mathcal{C}$ be a $3$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$ such that $\sum_{i=1}^{r}n_{i}=n$, and let $k$ denote the number of loose cycles with odd lengths. If $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq(n+2k)/4$, then $\mathcal{H}$ contains $\mathcal{C}$ as a spanning sub-hypergraph.
  • We also pose the following general problem concerning the embedding of disjoint loose cycles in a $k$-graph:
    Question 5.2.
    Let $k>3$ be a positive integer $\mathcal{C}$ be a $k$-graph consisting of vertex-disjoint loose cycles $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$ such that $\sum_{i=1}^{r}n_{i}=n$. Let $\mathcal{H}$ be an $n$-vertex $k$-graph. What is the minimum codegree condition on $\mathcal{H}$ that ensures $\mathcal{C}$ is a spanning sub-hypergraph of $\mathcal{H}$ when $n$ is sufficiently large?
  • Additionally, we consider the $1$-expansion version of the Hajnal-Szemerédi theorem, which leads to the following question:
    Question 5.3.
    Let $t$ be a positive integer and let $K^{+}_{t}$ denote the $1$-expansion of $K_{t}$. Given an $n$-vertex $3$-graph $\mathcal{H}$, what is the minimum codegree condition on $\mathcal{H}$ that ensures the existence of a $K^{+}_{t}$-factor when $n$ is sufficiently large and $t+t(t-1)/2$ divides $n$?
6 Acknowledgements
Sincere thanks to Professor Jie Han for his insightful suggestions for this paper.
Appendix A Appendix
A.1
In this subsection, we deduce Theorem 1.4 from Theorem 4.1. The prove is similar with the work of Khan in [16].
First, we prove that every $n$-vertex $3$-graph with minimum codegree at least $(1/4+o(1))n$ contains a loose cycle on any even number of vertices. For this, we need the following two results.
Lemma A.1 (Chvátal, 1979).
Let
$F(n,\lambda,k,\delta)=\binom{n}{k}^{-1}\sum_{i=0}^{(\lambda-\delta)k}\binom{\lambda n}{i}\binom{(1-\lambda)n}{k-i}$
denote the probability that a random subset of $k$ elements out of $n$, $\lambda n$ of which are marked, contains at most $(\lambda-\delta)k$ marked elements; then
$F(n,\lambda,k,\delta)\leq e^{-2\delta^{2}k}.$
Theorem A.2 ([9]).
There is an integer $n_{0}$ such that every $3$-graph $\mathcal{H}$ on $n\geq n_{0}$ vertices where $n$ is even and with minimum codegree at least $n/4$ contains a loose Hamilton cycle.
Lemma A.3.
For any $\gamma>0$, there exists a constant $N=N(\gamma)$ such that if $\mathcal{H}$ is a $3$-graph on $n\geq N$ vertices with minimum codegree $\delta_{2}(\mathcal{H})\geq(1/4+\gamma)n$, then $\mathcal{H}$ is super-pancyclic, i.e., $\mathcal{H}$ contains all loose cycles on $q$ vertices, where $q$ is an even integer and $4\leq q\leq n$.
Proof.
The proof is divided into two cases depending on $q$.
For $44\leq q\leq n/$ 4, choose a loose path $P=v_{1}v_{2}\ldots v_{q-1}$ on $(q-1)$ vertices (such a path exists as $\mathcal{H}$ contains a loose Hamilton cycle by Theorem 1.2). Since $\delta_{2}(\mathcal{H})\geq(1/4+\gamma)n$, we have $|N_{V(\mathcal{H})\backslash V(P)}(v_{1},v_{q-1})|\geq\gamma n>0$. Then, we can choose a vertex $v_{q}\in N_{V(\mathcal{H})\setminus V(P)}(v_{1},v_{q-1})$ such that $v_{1}v_{2}\ldots v_{q-1}v_{q}$ is a loose cycle on $q$ vertices, as desired.
For $q>n/4$, choose a subset $A\subseteq V(\mathcal{H})$ with $|A|=q$ among all such sets uniformly at random. For $u,v\in V(\mathcal{H})$, let $X_{u,v}$ be the random variable $|N(u,v)\cap A|$. Then, by Lemma A.1, we have
$\mathds{P}(X_{u,v}\leq q/4)\leq e^{-2\gamma^{2}q}\leq e^{-\gamma^{2}n/2}.$
By the union bound,
$\mathds{P}(\exists u,v\in V(\mathcal{H}):X_{u,v}\leq q/4)\leq n^{2}\max_{u,v\in V(\mathcal{H})}\mathds{P}(X_{u,v}\leq q/4)\leq n^{2}e^{-\gamma^{2}n/2}.$
We can choose $N$ such that for $n\geq N$, this probability tends to 0. This implies that there exists a subset $A\subseteq V(\mathcal{H})$ with $|A|=q>n/4$ such that for each pair $u,v\in A$, we have $X_{u,v}\geq|A|/4$. Therefore, there is a loose Hamilton cycle in $\mathcal{H}[A]$ by Theorem A.2. ∎
Recall that $C_{t}$ denotes a loose cycle on $t$ vertices. Suppose that $\mathcal{H}$ is an $n$-vertex $3$-graph with $\delta_{2}(\mathcal{H})\geq(n+2k)/4+\eta n\geq(1/4+\eta)n$, and $\mathcal{C}$ is a $3$-graph consisting of vertex-disjoint union of $C_{n_{1}},C_{n_{2}},\ldots,C_{n_{r}}$ satisfying $\sum_{i=1}^{r}n_{i}=n$. Let $\mathcal{C}_{1}$ be the set of loose cycles on at most $\log n$ vertices in $\mathcal{C}$ and $\mathcal{C}_{2}=\mathcal{C}\backslash\mathcal{C}_{1}$. Define $|V(\mathcal{C}_{1})|=cn$. We consider the following cases.
Case 1.
$20\leq c<\eta/$ 2 .
Using Lemma A.3, for given $\eta>0$, we can choose a suitable integer $N$ such that the loose cycles in $\mathcal{C}_{1}$ are embedded into $\mathcal{H}$ one by one. Indeed, let $\mathcal{H}^{\prime}=\mathcal{H}[V(\mathcal{H})\setminus V(\mathcal{C}_{1})]$, we have $\delta_{2}(\mathcal{H}^{\prime})\geq(1/4+\eta)n-\eta n/2\geq(1/4+\eta/2)|V(\mathcal{H}^{\prime})|$. We embed the loose cycles of $\mathcal{C}_{2}$ into $\mathcal{H}^{\prime}$ as Case 4.
Case 2.
$\eta/2\leq c<1-\eta/2.$
In this case, we randomly partition $V(\mathcal{H})$ into $V_{1}$ and $V_{2}$ with $|V_{1}|=|V(\mathcal{C}_{1})|$ and $|V_{2}|=|V(\mathcal{C}_{2})|$. For $u,v\in V(\mathcal{H})$, let $X_{u,v}^{i}$ be the random variable $|N_{\mathcal{H}}(u,v)\cap V_{i}|$, where $i\in[2]$. By Lemma A.1, we have
$\mathds{P}(X_{u,v}^{i}\leq(1/4+\eta/2)|V_{i}|)\leq e^{-\eta^{2}|V_{i}|/2}.$
By the union bound,
$\mathds{P}(\exists u_{i},v_{i}\in V_{i}:X_{u_{i},v_{i}}^{i}\leq(1/4+\eta/2)|V_{i}|)$ $\leq|V_{i}|^{2}\max_{u_{i},v_{i}\in V_{i}}\mathds{P}(X_{u_{i},v_{i}}^{i}\leq(1/4+\eta/2)|V_{i}|)$
$\leq|V_{i}|^{2}e^{-\eta^{2}|V_{i}|/2}.$
Therefore, we can choose $n_{0}$ such that for $n\geq n_{0}$, this probability tends to 0. This implies that there is a partition $V=V_{1}\cup V_{2}$ such that $\delta_{2}(\mathcal{H}[V_{1}])\geq(1/4+\eta/2)|V_{1}|$ and $\delta_{2}(\mathcal{H}[V_{2}])\geq(1/4+\eta/2)|V_{2}|$. Then, the loose cycles of $\mathcal{C}_{1}$ can be embedded into $\mathcal{H}[V_{1}]$ by Theorem 4.1, and the loose cycles of $\mathcal{C}_{2}$ can be embedded into $\mathcal{H}[V_{2}]$ as Case 4.
Case 3.
$11-\eta/2\leq c<$ 1 .
Note that $20<|V(\mathcal{C}_{2})|\leq\eta n/$ 2. We can embed loose cycles in $\mathcal{C}_{2}$ one by one into $\mathcal{H}$ by using Lemma A.3. Let $\mathcal{H}^{\prime}=\mathcal{H}[V\setminus V(\mathcal{C}_{2})]$, we know that $\delta_{2}(\mathcal{H}^{\prime})\geq(1/4+\eta/2)n\geq(1/4+\eta/2)|V(\mathcal{H}^{\prime})|$. Then the loose cycles of $\mathcal{C}_{1}$ can be embedded into $\mathcal{H}^{\prime}$ by Theorem 4.1.
Case 4.
$c=1$, i.e., all loose cycles in $\mathcal{C}$ has larger than $\log n$ vertices.
Assume that $n_{1}\geq n_{2}\geq\ldots\geq n_{r}>\log n$. In this case, we consider two cases relying on $n_{1}$.
Subcase 4.1 $n_{1}\geq(1-\eta/2)n.$
We embed all loose cycles except for $C_{n_{1}}$ one by one using Lemma A.3. Indeed, let $\mathcal{H}^{\prime}=\mathcal{H}[V(\mathcal{H})\setminus\cup_{i=2}^{r}V(C_{n_{i}})]$, we know that $\delta_{2}(\mathcal{H}^{\prime})\geq(1/4+\eta/2)n\geq(1/4+\eta/2)|V(\mathcal{H}^{\prime})|$. Then $\mathcal{H}^{\prime}$ contains a loose Hamilton cycle. Then, we embed all loose cycles.
Subcase 4.2 $n_{i}<(1-\eta/2)n$ for each $i\in[r].$
Divide the loose cycles of $\mathcal{C}$ into two parts $\mathcal{C}_{1}^{1}$ and $\mathcal{C}_{1}^{2}$ such that $|V(\mathcal{C}_{1}^{1})|\leq(1-\eta/2)n$ and $|V(\mathcal{C}_{1}^{2})|\leq(1-\eta/2)n$. Then we randomly partition $V(\mathcal{H})$ into two sets $V_{1}^{1}$ and $V_{1}^{2}$ such that $|V_{1}^{1}|=|V(\mathcal{C}_{1}^{1})|$ and $|V_{1}^{2}|=|V(\mathcal{C}_{1}^{2})|$. In the same way as Case 2, we know that there is a vertex partition $V=V_{1}^{1}\cup V_{1}^{2}$ such that $\delta_{2}(\mathcal{H}[V_{1}^{1}])\geq(1/4+\eta/2)|V_{1}^{1}|$ and $\delta_{2}(\mathcal{H}[V_{1}^{2}])\geq(1/4+\eta/2)|V_{1}^{2}|$. In the following, we will embed the loose cycles in $\mathcal{C}_{1}^{j}$ into $\mathcal{H}[V_{1}^{j}]$, respectively. If $\mathcal{C}_{1}^{j}$ contains a loose cycle on at least $(1-\eta/4)|V_{1}^{j}|$ vertices for some $j$, then $\mathcal{C}_{1}^{j}$ can be embedded into $\mathcal{H}[V_{1}^{j}]$ as Subcase 4.1. Otherwise, recursively apply the same partitioning process until the condition of Subcase 4.1 is satisfied. Let $k$ denote the times of the vertex partitions. Since the partitioning process terminates before the number of vertices in the sub-hypergraph is less than $\log n$, we have $k<n/\log n$. We will show that the minimum codegree condition holds till the end of the partitioning process. Indeed, by Lemma A.1 and the union bound, we have
$\mathds{P}\left(\exists u,v\in V_{k}^{j}:|N_{u,v}\cap V_{k}^{j}|\leq(1/4+\eta-\sum_{t=1}^{k}(\eta/2)^{t})|V_{k}^{j}|\right)$
$\leq|V_{k}^{j}|^{2}\max_{u,v\in V_{k}^{j}}\mathds{P}\left(|N_{u,v}\cap V_{k}^{j}|\leq(1/4+(1/2)^{k}\eta)|V_{k}^{j}|\right)$
$\leq|V_{k}^{j}|^{2}e^{-2(\eta-(1/2)^{k}\eta)^{2}|V_{k}^{j}|}\longrightarrow 0,$
where $j\in[2^{k}]$.
Hence, the minimum codegree conditions are satisfied in each partitioning process. We can apply the procedure of Subcase 4.1.
A.2
Lemma A.4.
For any positive integer $q$ and any sequence of positive numbers with $\sum_{i=1}^{k}a_{k}=1$, there exists a sequence of positive integers $q_{1},q_{2},\dots,q_{k}$ with $\sum_{i=1}^{k}q_{k}=q$ such that for any $i\in[k]$,
$|q_{i}-qa_{i}|\leq 1.$
Proof.
Let $q^{\prime}_{i}=\lfloor qa_{i}\rfloor$ for any $i\in[K].$ Let $w=q-\sum_{i=1}^{k}q^{\prime}_{i}$, we have:
$w\leq q-\sum_{i=1}^{k}(qa_{i}-1)\leq k.$
For every $i\in[w]$ let $q_{i}=q^{\prime}_{i}+1$ and for every $i\in[k]/[w]$ let $q_{i}=q^{\prime}_{i}.$ Thus, $\sum_{i=1}^{k}q_{i}=w+\sum_{i=1}^{k}q^{\prime}_{i}=q$, and for any $i\in[k]$,
$|q_{i}-qa_{i}|\leq\max\{qa_{i}-\lfloor qa_{i}\rfloor,\lceil qa_{i}\rceil-qa_{i}\}\leq 1.$